### v7

# Movement-Prediction-Adjusted Naïve Forecast

Abstract
This study introduces a movement-prediction-adjusted naïve forecast, which is the original naïve forecast with the addition of a weighted movement prediction term, in the context of forecasting time series that exhibit symmetric random walk properties. The weight of the movement term is determined by two parameters: one reflecting the directional accuracy and the other representing the mean absolute increment. The settings of the two parameters involve a trade-off: larger values may yield meaningful gains over the original naïve forecast, whereas smaller values often render the adjusted forecast more reliable. This trade-off can be managed by empirically setting the parameters using sliding windows on in-sample data. To statistically test the performance of the adjusted naïve forecast under different directional accuracy levels, we used four synthetic time series to simulate multiple forecast scenarios, assuming that for each directional accuracy level, diverse movement predictions were provided. The simulation results show that as the directional accuracy increases, the error of the adjusted naïve forecast decreases. In particular, the adjusted naïve forecast achieves statistically significant improvements over the original naïve forecast, even under a low directional accuracy of slightly above 0.50. This finding implies that the movement-prediction-adjusted naïve forecast can serve as a new optimal point forecast for time series with symmetric random walk characteristics if consistent movement prediction can be provided.
Keywords Naïve forecast  $\cdot$ Movement prediction  $\cdot$ Point forecasting  $\cdot$ Symmetric random walk
1 Introduction
Time series forecasting plays a crucial role in many decision-making processes. The widespread demand for accurate point forecasts across various domains has attracted considerable attention from researchers and practitioners. However, not all time series are equally predictable. Time series exhibiting random walk characteristics pose significant forecasting challenges (Hamilton,, 2020). Therefore, financial time series, known for these characteristics and their practical importance in financial markets, have been extensively studied as test cases for forecasting theories and methods. One prominent theory is the efficient market hypothesis (EMH), which suggests that asset prices incorporate all available information, rendering historical prices unpredictable (Fama,, 1970). For a random walk-like series, the naïve forecast, which assumes that the future value at the next time step is equal to the current value, is often regarded as the optimal point forecast.
In support of the EMH, Moosa, (2013) reported that as the volatility of financial time series increases, the root mean square error (RMSE) of the forecasting models increases more quickly than that of the naïve forecast. Moreover, Moosa and Burns, (2014) argued that claims of outperforming the naïve forecast in terms of the RMSE are often misleading, frequently involving introduced dynamics without rigorous statistical validation. Despite attempts to use various predictors to break the assumed randomness of financial time series, existing forecasting methods, including statistical, machine learning, and deep learning approaches, typically cannot outperform naïve forecasts (Kilian and Taylor,, 2003; Moosa and Burns,, 2016; Thakkar and Chaudhari,, 2021; Petropoulos et al.,, 2022; Ellwanger and Snudden,, 2023; Hewamalage et al.,, 2023; Zeng et al.,, 2023; Beck et al.,, 2025).
From the forecaster’s perspective, the unpredictability of random walk-like series could be attributed to incomplete knowledge of the underlying processes controlling the data, ignorance of the various contributing factors, and imperfections in analytical methods (Laplace,, 2012). In other words, the randomness of data is not inherent but rather apparent (Crutchfield and Feldman,, 2003). Although useful information may be contained in the input space, it can be overwhelmed by irrelevant data, making it difficult for current regression methods to exploit such information (John et al.,, 1994; Guyon and Elisseeff,, 2003; Verleysen and François,, 2005). Moreover, identifying the predictors on which the target variable truly depends is a complex task (Granger,, 1969; Stock and Watson,, 2002). However, these theories and empirical evidence reveal only the practical difficulty of finding exploitable patterns; they do not eliminate the possibility of discovering any predictability in the data.
Given the challenge in the point forecasting of random walk-like series, many researchers have shifted their focus from point forecasting to movement prediction, which involves predicting whether the value is increasing or decreasing for the next time step. Consequently, the regression task was reduced to a binary classification task. For a pure random walk, the probability that future movement will increase or decrease is 0.5, and the mutual information between the predictors and the target is zero (Kraskov et al.,, 2004). If the mutual information is greater than zero, it indicates that there is some dependency between the predictors and the target (Cover and Thomas,, 2012). Practically, this means that the predictors reduce the uncertainty of the binary outcome. Although this does not guarantee high predictive accuracy, it suggests that, theoretically, one can build a model that performs better than random guessing (Ircio et al.,, 2020). In contrast, for point forecasting in a random walk-like series, the continuous nature of the target means that the small gain in mutual information only marginally reduces the overall uncertainty and does not sufficiently narrow down the precise numerical result (Beraha et al.,, 2019).
Movement prediction is often considered easier than point forecasting in financial time series (Taylor,, 2008; Moosa,, 2013). Studies in the field of behavioral finance suggest that investor sentiment and psychology can lead to certain patterns (De Bondt and Thaler,, 1985; Lee et al.,, 1991; Baker and Wurgler,, 2007; Barberis et al.,, 1998), which could lead to mutual information between the sentiment predictors and the binary target being greater than zero and making future movements predictable. Therefore, sentiment indicators are often used to improve the directional accuracy in these tasks. For example, Weng et al., (2017) incorporated Google news counts and Wikipedia page views alongside historical data to predict the next day’s movement of AAPL stock via support vector machines (SVMs), achieving up to 0.80 directional accuracy. Ma et al., (2023) proposed a multisource aggregated classification (MAC) method for the prediction of stock price movement, incorporating the numerical features and market-driven news sentiments of target stocks, as well as the news sentiments of their related stocks, and the directional accuracy can reach as high as 0.70. Wang et al., (2025) used multimodal data that included news data to train a hybrid model consisting of a long short memory network (LSTM) and transformer for the prediction of stock movement, and the directional accuracy fell within the range of 0.57–0.87. Depending on the complexity of the markets and availability of data, the directional accuracy of financial time series typically ranges between 0.55 and 0.80 (Bustos and Pomares-Quimbaya,, 2020).
When a movement prediction is better than a random guess, it also indicates that the input data contain useful information that has been leveraged effectively to predict future movements. Borrowing the "less is more" principle from cognitive science, where reducing information complexity can lead to an improved cognitive process (Gigerenzer and Brighton,, 2009), we can subsequently use the result of the movement prediction to generate the point forecast, using only the binary labels as an additional predictor along with the original naïve forecast, as shown in Figure 1. The binary labels can then be regarded as "verified predictors" for the target variable, and no more noise is introduced to affect the final point forecast. Unfortunately, the use of movement prediction results only to improve point forecasts remains unexplored in the forecasting community. Most studies either stopped at the movement prediction stage or used a direct regression approach to conduct point forecasting. Although it is often argued that, in many cases, compared with point forecasting, movement prediction is more important for decision-making (Sezer et al.,, 2020), the final decision generally still depends on the estimated value of the target variable. For example, in financial markets, the decision could be to trade at a specific price given a predicted future movement. Therefore, the decision maker actually conducted the classification-to-regression conversion, albeit implicitly. However, how and why movement prediction results can be systematically exploited to improve point forecasting in random walk-like series is largely unknown.
Figure 1: Direct and indirect regression approaches
Under these conditions, this study aimed to explore the potential of using movement prediction to enhance point forecasting in random walk-like time series. In particular, the target variable is assumed to follow a symmetric random walk process, which is the basic type of random walk characterized by the absence of drift and equal probabilities of upward and downward movements (Ibe,, 2014). By reformulating the symmetric random walk process to incorporate the future movement of the target variable, the point forecast can naturally be expressed as the original naïve forecast adjusted by movement prediction, which is referred to as the movement-prediction-adjusted naïve forecast. The weight applied to the movement term is determined by two parameters: one reflecting the directional accuracy and the other representing the mean absolute increment. The setting of these two parameters introduces an inherent trade-off and uncertainty into the prediction performance. Specifically, larger parameter values suggest potentially meaningful improvements over the original naïve forecast, whereas smaller parameter values indicate a greater likelihood that the adjusted naïve forecast will outperform the original naïve forecast. To balance this trade-off, the parameters can be empirically determined by sliding windows on the in-sample data.
Since, in practice, movement prediction can exhibit a wide range of directional accuracies, the adjusted naïve forecast should be evaluated across different accuracy levels to examine the impact of directional accuracy on the performance of the proposed method. Therefore, we use synthetic time series to simulate multiple forecasting scenarios, assuming that for each directional accuracy level, diverse movement predictions are provided. For all the experiments, the evaluation metrics include the RMSE, mean absolute error (MAE), mean absolute percentage error (MAPE), and symmetric mean absolute percentage error (sMAPE), and the original naïve forecast serves as the sole baseline. In this experimental setting, it can be determined whether the improvement over the baseline by the proposed method is statistically significant.
This study makes three significant contributions to the literature. First, this study provides new insights into the predictability of random walk-like series, addressing a longstanding challenge in the field. When meaningful movement prediction was provided, the point values of the random walk-like series became partially predictable. Second, by illuminating the relationship between the actual and estimated parameter values, this study underscores the source of uncertainty in the effectiveness of the proposed method, a critical consideration often underemphasized by purely data-driven approaches. Third, the proposed method bridges movement prediction and point forecasting, which are two distinct tasks in time series forecasting. Although the use of binary classification to guide regression is well established, relying solely on the movement prediction of the target variable for its own point forecasting offers a fresh perspective.
The remainder of this paper is organized as follows. Section 2 introduces the details of the movement-prediction-adjusted naïve forecast, explains how to derive it, and explains why it can outperform the original naïve forecast. Section 3 presents a simulation study that assesses the performance of movement-prediction-adjusted naïve forecasts using synthetic datasets in a controlled experimental setting. Section 4 discusses the key findings, implications, and limitations of the study. Finally, Section 5 concludes the paper and outlines future research directions.
2 Movement-Prediction-Adjusted Naïve Forecast
This section presents the details of the movement-prediction-adjusted naïve forecast for time series that exhibit symmetric random walk characteristics. We first derive the general expression of the adjusted naïve forecast based on the definition of symmetric random walks, then explain how to optimize the parameters attached to the movement term, and finally, how to adjust the parameters to improve the forecasting effectiveness and detail the procedure for generating forecasts on the out-of-sample set.
2.1 General Expression
A symmetric random walk is a stochastic process in which the value at time step $t$ is the sum of the value at the previous time step $t-1$ and the value of an independent random increment ${{\epsilon}_{t}}$, which is generally assumed to follow a normal distribution with a mean of 0 and variance of $\sigma_{t}^{2}$ (De Gooijer et al.,, 2017). The formulation of symmetric random walks is as follows:
${{y}_{t}}={{y}_{t-1}}+{{\epsilon}_{t}},{{\epsilon}_{t}}\sim\mathcal{N}(0,\sigma_{t}^{2}),$ (1)
where ${y}_{t}$ is the point value at time step $t$ and ${y}_{t-1}$ is the point value at time step $t-1$. For random walks exhibiting homoscedasticity, the variance $\sigma_{t}^{2}$ remains constant over time, whereas for those exhibiting heteroscedasticity, the variance $\sigma_{t}^{2}$ varies over time. The symmetric nature of the random increment ${{\epsilon}_{t}}$ with a mean of 0 ensures that there is no inherent bias in the movement direction of the target variable and that upward and downward movements are equally likely (Ibe,, 2014). Furthermore, the random increment $\epsilon_{t}$ can be expressed as a product of two components: a nonnegative random increase $|\epsilon_{t}|$ representing the magnitude of the increment and a sign component ${D}_{t}$, which takes a value of $+1$ for upward movement and $-1$ for downward movement, indicating the direction of the increment between two adjacent points. This allows the process to be reformulated as follows:
${{y}_{t}}={{y}_{t-1}}+{{D}_{t}}\cdot|\epsilon_{t}|;|\epsilon_{t}|\geq 0$ (2)
The nonnegative random increment $|\epsilon_{t}|$ can be further expressed as a product of two terms:
${{y}_{t}}={{y}_{t-1}}+{{D}_{t}}\cdot{\theta_{t}}\cdot\bar{\epsilon};{\theta_{t}}\geq 0,\bar{\epsilon}>0,$ (3)
where $\bar{\epsilon}$ represents the mean absolute increment calculated from the in-sample data, which can be interpreted as a base magnitude of the increments. The term $\theta_{t}$ is introduced as a nonnegative parameter that controls the magnitude of the increment at time step $t$. Eq. (3) implies that the point forecast of the target variable can be generated if ${{D}_{t}}$ can be predicted and both $\theta_{t}$ and $\bar{\epsilon}$ can be estimated. Consider a simple scenario in which the magnitude of the increment between two adjacent points is fixed for future time steps; then, $\theta_{t}$ can be set to a constant value. Consequently, the point forecast for future time step $t$, denoted as $\hat{y}_{t}$, can be expressed as follows:
${{\hat{y}}_{t}}={{y}_{t-1}}+\hat{D}_{t}\cdot\theta\cdot\bar{\epsilon};\bar{\epsilon}>0,$ (4)
where ${{\hat{D}}_{t}}$ is the movement prediction, $\theta$ is a fixed scalar, and $\theta\in\mathbb{Q}$. Therefore, the point forecast generated via Eq. (4) can be regarded as an adjustment of the original naïve forecast through a weighted movement prediction term; hence, it is named the movement-prediction-adjusted naïve forecast.
2.2 Movement Term Weight
According to Eq. (4), the weight of the movement term in the adjusted naïve forecast is the product of two parameters, $\theta$ and $\bar{\epsilon}$, both of which can be estimated from the in-sample data. Since $\bar{\epsilon}$ is directly computed as the mean absolute increment of the in-sample series, only $\theta$ requires optimization. This is typically achieved by minimizing the mean squared error (MSE) of the adjusted naïve forecast for the in-sample set. Because the MSE of the original naïve forecast is a fixed value on the in-sample set, minimizing the MSE of the adjusted naïve forecast is mathematically equivalent to maximizing the difference between the MSEs of the original and adjusted naïve forecasts:
$\min MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\equiv\max\left(MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\right)$ (5)
Using this approach, we can determine the optimal value of $\theta$ and examine, through the difference between the MSEs, whether there is an improvement over the original naïve forecast under this optimal setting.
Suppose that the in-sample set has $N$ time steps and that movement prediction $\hat{D}_{t}$ is provided for each time step. Within the in-sample set, we assume that the movement prediction yields $n_{\text{in-sample}}^{\text{correct}}$ correct and $n_{\text{in-sample}}^{\text{incorrect}}$ incorrect directions. We subsequently used Eq. (2) to present the actual value and expand the MSE expression for the original naïve forecast:
$MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (6)
$=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t-1}+D_{t}\cdot|\epsilon_{t}|-y_{t-1}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\ |\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}\ |\epsilon_{t}|^{2}\right)$
When using Eq. (2) to represent the actual value and Eq. (4) to represent the predicted value, the MSE of the adjusted naïve forecast on the in-sample set can be expressed as:
$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$ (7)
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(y_{t}-\hat{y}_{t}\right)^{2}+\sum_{t\in\text{incorrect}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left[(y_{t-1}+D_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{D}_{t}\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}})\right]^{2}\right.$
$\quad\left.+\sum_{t\in\text{incorrect}}\left[(y_{t-1}+D_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{D}_{t}\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}})\right]^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right),$
where $\theta_{\text{in-sample}}$ represents the parameter $\theta$ estimated from the in-sample set and $\bar{\epsilon}_{\text{in-sample}}$ denotes the mean of $|\epsilon_{t}|$ calculated over the in-sample set. Consequently, the difference between the two MSEs on the in-sample set can be derived via Eq. (6) and Eq. (7), respectively:
$\Delta\text{MSE}_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (8)
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
$\quad-\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
$\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
Based on the law of large numbers (LLN), when both $n_{\text{in-sample}}^{\text{correct}}$ and $n_{\text{in-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}}$ (9)
Therefore, Eq. (8) can be simplified as follows:
$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (10)
$\approx\frac{1}{N}\left(2\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}})-n_{\text{in-sample}}^{\text{correct}}\cdot\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right.$
$\quad\left.-2\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}})-n_{\text{in-sample}}^{\text{incorrect}}\cdot\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)$
$=(4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}})\cdot\bar{\epsilon}_{\text{in-sample}}^{2},$
where $ACC_{\text{in-sample}}$ is the directional accuracy of the movement prediction on the in-sample set, which is defined as:
$ACC_{\text{in-sample}}=\frac{n_{\text{in-sample}}^{\text{correct}}}{n_{\text{in-sample}}^{\text{incorrect}}+n_{\text{in-sample}}^{\text{correct}}}$ (11)
To maximize the $\Delta MSE_{\text{in-sample}}$, we focus on the expression inside parentheses in the final form of Eq. (10), which can be regarded as a function of parameter $\theta_{\text{in-sample}}$:
$f(\theta_{\text{in-sample}})=4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}}$ (12)
By differentiating $f(\theta_{\text{in-sample}})$ and setting $f^{\prime}(\theta_{\text{in-sample}})=0$, the critical value of $\theta_{\text{in-sample}}$, denoted as $\theta_{\text{in-sample}}^{*}$, is given by:
$\theta_{\text{in-sample}}^{*}=2\cdot ACC_{\text{in-sample}}-1$ (13)
Because $f^{\prime\prime}(\theta_{\text{in-sample}})<0$, the function $f(\theta_{\text{in-sample}})$ is concave downward, and the critical point of $f(\theta_{\text{in-sample}})$ is the maximum. Therefore, $\theta_{\text{in-sample}}^{*}$ represents the optimal value of parameter $\theta$ in the in-sample set. In addition, Eq. (13) reveals the relationship between the optimal value of parameter $\theta$ and the directional accuracy of the movement prediction within a fixed sequence length. According to Eqs. (10) and (13), the improvement over the original naïve forecast on the in-sample set, achieved by the adjusted naïve forecast with optimal parameters, is expressed as follows:
$\max\Delta MSE_{\text{in-sample}}$ $=\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$ (14)
According to Eqs. (13) and (14), as the directional accuracy increases from 0.5, the value of $f(\theta_{\text{in-sample}})$ increases from 0, and the improvement over the original naïve forecast correspondingly increases from 0. Because this improvement is a nonnegative value, the adjusted naïve forecast is at least as good as the original naïve forecast on the in-sample set:
$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\leq MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ (15)
2.3 Forecast on the Out-of-Sample Set
If we use $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as the estimates of the actual optimal values of $\theta$ and $\bar{\epsilon}$ on the out-of-sample set and suppose that the movement prediction on the out-of-sample set with $M$ time steps yields $m_{\text{out-of-sample}}^{\text{correct}}$ correct directions and $m_{\text{out-of-sample}}^{\text{incorrect}}$ incorrect directions; then, the difference between the two MSEs on the out-of-sample set can be expressed as follows:
$\Delta\text{MSE}_{\text{out-of-sample}}$ $=\text{MSE}_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (16)
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
$\quad-\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
$\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
Let $\bar{\epsilon}_{\text{out-of-sample}}$ denote the actual mean absolute increment on the out-of-sample set. According to the LLN, when both $m_{\text{out-of-sample}}^{\text{correct}}$ and $m_{\text{out-of-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}}$ (17)
Therefore, Eq. (16) can further be expressed as:
$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (18)
$\approx\frac{1}{M}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}})-m_{\text{out-of-sample}}^{\text{correct}}\cdot\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right.$
$\quad\left.-2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}})-m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)$
$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot ACC_{\text{out-of-sample}}\cdot\bar{\epsilon}_{\text{out-of-sample}}-2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}$
$\quad\cdot(1-ACC_{\text{out-of-sample}})\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2},$
where $ACC_{\text{out-of-sample}}$ is the directional accuracy of the movement prediction on the out-of-sample set, which is defined as:
$ACC_{\text{out-of-sample}}=\frac{m_{\text{out-of-sample}}^{\text{correct}}}{m_{\text{out-of-sample}}^{\text{incorrect}}+m_{\text{out-of-sample}}^{\text{correct}}},$ (19)
and $\theta_{\text{out-of-sample}}^{*}$ is the actual optimal value of $\theta$ for the out-of-sample set, which is derived according to Eq. (13):
$\theta_{\text{out-of-sample}}^{*}=2\cdot ACC_{\text{out-of-sample}}-1$ (20)
To ensure the effectiveness of the adjusted naïve forecast, the $\Delta MSE_{\text{out-of-sample}}$ should be no less than zero. Therefore, according to Eq. (18), the following inequality must be satisfied:
$\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}\geq\frac{1}{2}\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}$ (21)
Given that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are specific estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, respectively, inequality (21) can be generalized as:
$\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}\geq\frac{1}{2}\cdot\hat{\theta}_{\text{out-of-sample}}\cdot\hat{\bar{\epsilon}}_{\text{out-of-sample}},$ (22)
where $\hat{\theta}_{\text{out-of-sample}}$ represents the estimate of $\theta_{\text{out-of-sample}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ represents the estimate of $\bar{\epsilon}_{\text{out-of-sample}}$. Since the values of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ cannot be determined in advance, whether the inequality (22) can be satisfied is unknown, even when the optimal parameters derived from the in-sample set are used. Therefore, inequality (22) reveals the source of the uncertainty in the effectiveness of the adjusted naïve forecast, which depends on whether the product of the actual parameters exceeds half the product of the estimated parameters.
We can regard $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ as two variables, each of which follows a distribution with a nonnegative range. Inequality (22) represents an event whose probability of occurrence depends on the estimated values of the two variables. When both estimates are set to small values near zero, the probability that inequality (22) is satisfied approaches one. However, in such a setting, the improvement over the original naïve forecast by the adjusted naïve forecast on the out-of-sample set also diminishes. Therefore, inequality (22) also reflects a trade-off between maximizing the probability that the adjusted naïve forecast outperforms the original forecast and achieving a meaningful level of improvement.
In addition to using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, which may result in overfitting, a more conservative approach can be used to set the estimated values of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$:
  1. 1.
    Consider a data-generating process in which a sliding window, whose length is equal to that of the out-of-sample set, moves sequentially from the beginning to the end of the in-sample set. Each window can be treated as an "out-of-sample set" in the historical context.
  2. 2.
    Within each window, the directional accuracy is calculated and then translated into the corresponding value of $\theta$ via Eq. (13). As the window moves, a collection of $\theta$ values is generated, denoted as $\Theta_{\text{set}}$, whose bounds can be regarded as the estimated range of $\theta_{\text{out-of-sample}}^{*}$.
  3. 3.
    Similarly, a sliding window of the same length is applied to calculate a set of $\bar{\epsilon}$ values from the in-sample set, denoted as $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, providing an estimated value range of $\bar{\epsilon}_{\text{out-of-sample}}$.
  4. 4.
    Using $\min(\Theta_{\text{set}})$ and $\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})$ as conservative estimates for $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, the adjusted naïve forecast for time step $t$ in the out-of-sample set can be expressed as:
    $\hat{y}_{t}=y_{t-1}+\hat{D}_{t}\cdot\min(\Theta_{\text{set}})\cdot\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})$ (23)
Algorithm 1 describes the process of generating the adjusted naïve forecast on the entire out-of-sample set.
Algorithm 1 Movement-Prediction-Adjusted Naïve Forecast on the Out-of-sample Set
Inputs:
$Y_{\text{in-sample}}$ //    Array of observations in the in-sample set
$Y_{\text{out-of-sample}}$ //    Array of observations in the out-of-sample set
$\hat{D}_{\text{in-sample}}$ //    Movement prediction on the in-sample set
$\hat{D}_{\text{out-of-sample}}$ //    Movement prediction on the out-of-sample set
$N$ //    Number of elements in the in-sample set
$M$ //    Number of elements in the out-of-sample set
Output:
$P$ //    Adjusted naïve forecast on the out-of-sample set
1. Initialize variables:
$D_{\text{in-sample}}\leftarrow[]$ //    Actual movements on the in-sample set
$P\leftarrow[]$
$ACC_{\text{set}}\leftarrow[]$ //    Directional accuracy set
$\Theta_{\text{set}}\leftarrow[]$ //    Theta set
$R_{\text{in-sample}}\leftarrow[]$ //    First-order difference set
$\boldsymbol{\bar{\epsilon}}_{\text{set}}\leftarrow[]$ //    Mean absolute increment set
2. Calculate actual movements:
for $t\leftarrow 1$ to $\text{length}(Y_{\text{in-sample}})$ do
     $\text{diff}\leftarrow Y_{\text{in-sample}}[t]-Y_{\text{in-sample}}[t-1]$ //    First-order difference
     $R_{\text{in-sample}}.\text{append}(\text{diff})$
     if $\text{diff}>0$ then
         $D_{\text{in-sample}}.\text{append}(1)$ //    Upward movement
     else
         $D_{\text{in-sample}}.\text{append}(-1)$ //    Downward or no movement
     end if
end for
3. Calculate directional accuracies:
for $t\leftarrow 1$ to $N-M+1$ do
[…]
     for $i\leftarrow t$ to $t+M-1$ do
         if $D_{\text{in-sample}}[i]==\hat{D}_{\text{in-sample}}[i]$ then
[…]
         end if
     end for
     $ACC_{\text{set}}.\text{append}(\text{Correct\_Predictions}/M)$
end for
4. Compute the movement weights:
for each $\text{each }ACC\in ACC_{\text{set}}$ do
     $\theta\leftarrow 2\cdot\text{ACC}-1$
     $\Theta_{\text{set}}.\text{append}(\theta)$
end for
5. Compute the mean absolute increments:
for $t\leftarrow 1$ to $N-M+1$ do
     $\bar{\epsilon}\leftarrow\text{mean}(|R_{\text{in-sample}}[t:t+M-1]|)$
     $\boldsymbol{\bar{\epsilon}}_{\text{set}}.\text{append}(\bar{\epsilon})$
end for
6. Set the initial forecast:
$Y_{\text{last}}\leftarrow Y_{\text{in-sample}}[\text{end}]$ //    Start from the last element of the in-sample set
7. Forecast the out-of-sample set:
for $i\leftarrow 1$ to $\text{length}(Y_{\text{out-of-sample}})-1$ do
     $\Delta\leftarrow\min(\Theta_{\text{set}})\cdot\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})\cdot\hat{D}_{\text{out-of-sample}}[i]$ //    Compute increment
     $Y_{\text{forecast}}\leftarrow Y_{\text{last}}+\Delta$ //    Compute the adjusted naïve forecast
     $P.\text{append}(Y_{\text{forecast}})$ //    Store result
     $Y_{\text{last}}\leftarrow Y_{\text{out-of-sample}}[i]$
end for
8. Return:
return $P$
3 Simulation
This section presents a simulation study using four synthetic time series to evaluate whether the adjusted naïve forecast outperforms the original naïve forecast under different directional accuracy levels; therefore, movement prediction is assumed to have been provided. To statistically test the performance of the proposed method on an out-of-sample set, multiple movement predictions were randomly generated for each directional accuracy level.
3.1 Datasets and Setup
The dataset includes four synthetic time series, each of which represents a symmetric random walk with a certain type of variance, as summarized in Table 1. Each time series, consisting of 2,500 steps, used a fixed random seed (seed = 1) in the Python environment, as shown in Figure 2. To maintain all values within the positive range, an initial positive offset of 10000 units was applied to each series. The last 500 data points were selected as the out-of-sample set, and the remaining data points were used as the in-sample set. In addition, Figure 3 shows the first-order differences of each time series to visually demonstrate the properties of the increments, and Figure 4 shows the histograms of the first-order differences for all time series, highlighting the symmetric nature of the time series.
Table 1: Synthetic time series summary
Variance Type Expression Parameter Range
Constant variance $\sigma_{t}^{2}=\sigma_{0}^{2}$ $\sigma_{0}^{2}=1$
Linear variance trend $\sigma_{t}^{2}=\frac{\sigma_{0}^{2}}{1+k\cdot t}$ $\sigma_{0}^{2}=1$, $k\in(0,10)$
Cyclic variance $\sigma_{t}^{2}=\sigma_{0}^{2}(1+a\cdot\sin(\omega t))$ $\sigma_{0}^{2}=1$, $\omega=\frac{2\pi}{100}$, $a\in(0,10)$
Randomly varying variance $\sigma_{t}^{2}=\sigma_{t-1}^{2}+\eta_{t},\eta_{t}\sim\mathcal{N}(0,\xi^{2})$ $\sigma_{0}^{2}=1$, $\xi^{2}\in(0,1000)$
Figure 2: Synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
Figure 3: First-order differences of four synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
Figure 4: Histograms of first-order differences of four synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
For each time series, the adjusted naïve forecast on the out-of-sample set was generated via Algorithm 1, which uses the values of $\theta$ determined by different levels of predefined directional accuracies. The performance of each adjusted naïve forecast result was evaluated using four metrics: the RMSE, MAE, MAPE, and sMAPE. The RMSE, derived from the square root of the MSE, emphasizes larger errors, which makes it suitable for identifying significant deviations in forecasts. The MAE, on the other hand, provides an interpretable measure of the average error without overpenalizing outliers. The MAPE and sMAPE normalize the errors relative to the actual values, enabling comparisons between datasets with different scales. Together, these metrics capture both absolute and relative error characteristics, ensuring a robust and balanced evaluation. The formulas for these metrics are as follows:
$RMSE=\sqrt{\frac{1}{N}\sum\limits_{i=1}^{N}{{{({{y}_{i}}-{{{\hat{y}}}_{i}})}^{2}}}},$ (24)
$MAE=\frac{1}{N}\sum\limits_{i=1}^{N}{\left|{{y}_{i}}-{{{\hat{y}}}_{i}}\right|},$ (25)
$MAPE=\frac{100}{N}\sum\limits_{i=1}^{N}{\left|\frac{{{y}_{i}}-{{{\hat{y}}}_{i}}}{{{y}_{i}}}\right|},$ (26)
$sMAPE=\frac{100}{N}\sum\limits_{i=1}^{N}{\frac{\left|{{y}_{i}}-{{{\hat{y}}}_{i}}\right|}{\frac{\left|{{y}_{i}}\right|+\left|{{{\hat{y}}}_{i}}\right|}{2}}},$ (27)
where $N$ denotes the size of the out-of-sample set, ${{\hat{y}}_{i}}$ denotes the predicted value, and ${{y}_{i}}$ denotes the actual value.
To statistically test the performance of the adjusted naïve forecast, for each synthetic time series at a specified directional accuracy level, the process to generate the adjusted naïve forecast was repeated 100 times. Each repetition used a unique movement prediction sequence at the same directional accuracy level, which was generated by randomly replacing a subset of actual movements in the out-of-sample set with incorrect movements. The directional accuracy level varies in two ways: initially, it changes from 0.50 to 0.56 in increments of 0.01, and then it changes from 0.55 to 1.00 in larger increments of 0.05. This approach allows for finer granularity in the lower accuracy range, where small changes in the directional accuracy can have a significant effect on the performance of the method, while using larger steps at higher accuracy levels balances the computational efficiency. Consequently, for each synthetic time series, 100 independent values were calculated for each metric at each accuracy level.
3.2 Results and Analysis
Because the directional accuracy for the out-of-sample set is predefined in each simulation process, only the value of $\bar{\epsilon}_{\text{out-of-sample}}$ needs to be estimated. Figure 5 illustrates the distribution of the values in the $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, generated on the basis of sliding windows across the first-order differences of the in-sample set. The vertical reference line in Figure 5 represents the actual $\bar{\epsilon}_{\text{out-of-sample}}$ value of the out-of-sample set. In the four cases, the minimum value of $\boldsymbol{\bar{\epsilon}}_{\text{set}}$ is generally an underestimate of $\bar{\epsilon}_{\text{out-of-sample}}$, except for the second case, where the synthetic time series is a random walk with a decreasing trend of linear variance. However, the estimated value remains close to the true value. Under the assumption that directional accuracy is consistent across the in-sample and out-of-sample sets, this estimation ensures that inequality (22) is satisfied.
Figure 5: Comparison of the true $\bar{\epsilon}_{\text{out-of-sample}}$ with $\boldsymbol{\bar{\epsilon}}_{\text{set}}$: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
Furthermore, comparisons of the two types of naïve forecasts based on the four evaluation metrics at low accuracy levels are illustrated in Figures 6, 7, 8, and 9. Each box plot in each figure represents the distribution of 100 independent metric values at a specific directional accuracy level, with the metric value of the original naïve forecast shown as a reference line. According to Figure 6, at low accuracy levels, such as 0.51, the RMSE of the adjusted naïve forecast is comparable to or slightly better than that of the original naïve forecast, as shown by the proximity of the median of each box plot to the reference line. As the accuracy increased beyond 0.52, most of the RMSE values fell below the baseline.
Figure 6: RMSEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 7: MAEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 8: MAPEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 9: sMAPEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Because the adjusted naïve forecast consistently outperforms the original naïve forecast at low accuracy levels in terms of the MAE, MAPE, and sMAPE, the Wilcoxon signed rank test was conducted to specifically evaluate whether the median differences between the RMSE values of the adjusted naïve forecast and those of the original naïve forecast significantly deviated from zero for accuracy levels ranging from 0.50–0.56 (Li and Johnson,, 2014). As summarized in Table 2, at a directional accuracy level of 0.51, the adjusted naïve forecast shows mixed results with respect to statistical significance. Although significant improvements were observed over the original naïve forecast for the Type 4 series (randomly varying variance, p = 0.002), the results for the Type 1 (constant variance, p = 0.915), Type 2 (linear variance trend, p = 0.781), and Type 3 (cyclic variance, p = 0.812) series were not statistically significant, indicating that an accuracy level of 0.51 may not consistently produce meaningful improvements across all types of time series. As the accuracy increased beyond 0.51, significant reductions in the RMSE (p < 0.001) were observed for all types of time series, thereby demonstrating the reliability and effectiveness of the adjusted naïve forecast at higher accuracy levels.
Table 2: RMSE results of the adjusted naïve forecast across directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Prediction Type 1 Type 2 Type 3 Type 4
RMSE Range p-value RMSE Range p-value RMSE Range p-value RMSE Range p-value
Original Naïve 0.961 - 0.00762 - 1.355 - 21.034 -
Acc = 0.50 0.961 - 0.00762 - 1.355 - 21.034 -
Acc = 0.51 0.960 $\pm$ 0.0004 0.915 0.00762 $\pm$ 0.000003 0.781 1.355 $\pm$ 0.0005 0.812 21.032 $\pm$ 0.0053 0.002
Acc = 0.52 0.960 $\pm$ 0.0008 <0.001 0.00762 $\pm$ 0.000007 <0.001 1.355 $\pm$ 0.0008 <0.001 21.024 $\pm$ 0.0102 <0.001
Acc = 0.53 0.959 $\pm$ 0.0012 <0.001 0.00761 $\pm$ 0.000010 <0.001 1.354 $\pm$ 0.0015 <0.001 21.015 $\pm$ 0.0144 <0.001
Acc = 0.54 0.959 $\pm$ 0.0017 <0.001 0.00761 $\pm$ 0.000014 <0.001 1.353 $\pm$ 0.0020 <0.001 20.995 $\pm$ 0.0199 <0.001
Acc = 0.55 0.957 $\pm$ 0.0020 <0.001 0.00760 $\pm$ 0.000020 <0.001 1.352 $\pm$ 0.0028 <0.001 20.979 $\pm$ 0.0244 <0.001
Acc = 0.56 0.956 $\pm$ 0.0023 <0.001 0.00758 $\pm$ 0.000021 <0.001 1.351 $\pm$ 0.0032 <0.001 20.957 $\pm$ 0.0278 <0.001
According to Figures 10, 11, 12, and 13, at higher accuracy levels ranging from 0.55–1.00, the adjusted naïve forecast consistently outperforms the original naïve forecast in terms of the four metrics. In particular, the improvement over the original naïve forecast becomes more pronounced as the directional accuracy improves. This finding demonstrates that highly accurate movement predictions can be effectively translated into improved point forecasts in the context of forecasting random walk-like time series. However, the improvement over the original naïve forecast is limited, because the proposed method allows a random walk-like series to be predicted to a certain extent, but the predicted value cannot be infinitely close to the actual value.
Figure 10: RMSEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 11: MAEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 12: MAPEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 13: sMAPEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
In general, when the directional accuracy was 0.5, the adjusted naïve forecast did not provide an improvement over the original naïve forecast. However, as the directional accuracy increases, the performance of the adjusted naïve forecast improves. This result is consistent with the theoretical analysis of the proposed method, which highlights its robustness and potential applicability to a wide range of time series that exhibit symmetric random walk characteristics, including those with homoscedasticity as well as high-volatility scenarios.
4 Discussion
This section discusses the implications of the findings and potential limitations of this study. First, this study demonstrates an indirect regression approach that uses only movement predictions and the original naïve forecast for a random walk-like time series. The theoretical analysis and simulation results indicate that as the directional accuracy increases, the improvement over the original naïve forecast also increases, suggesting that improving movement predictions may be more critical than developing direct regression models in this context. It also highlights a fundamental trade-off between maximizing the probability that the method outperforms the original naïve forecast and achieving a meaningful level of improvement, reflecting an inherent uncertainty in prediction tasks that purely data-driven approaches may sometimes overlook.
Second, many real-world forecasting scenarios require the integration of predictions from different data types to support decision-making (Bishop and Nasrabadi,, 2006; Timmermann,, 2006). However, it is difficult to combine binary and continuous forecasts. The proposed method can be regarded as a feasible approach to convert binary movement predictions into continuous point forecasts. This conversion facilitates the comprehensive use of diverse predictions and can enhance decision-making in practice.
Third, this study focuses on the theoretical predictability of random walk-like time series rather than exploiting this predictability for profit. Although certain structural properties or external signals may slightly improve movement predictions and thereby point forecasts, the forecasting result does not necessarily translate into economic gains, especially when transaction costs, market frictions, and slippage are considered.
Finally, because the aim of this study was to show how incorporating movement predictions can enhance naïve forecasts rather than how to generate movement predictions, we relied on simulations rather than real-world examples. Notably, the effectiveness of the adjusted naïve forecast depends on the consistent directional accuracy across different segments of the time series. Movement prediction models prone to overfitting may show a decline in directional accuracy on the out-of-sample set, limiting the actual improvement over the naïve forecast using the proposed method.
5 Conclusion
This study introduces a movement-prediction-adjusted naïve forecast, which is a linear combination of movement prediction and the original naïve forecast, in the challenging context of forecasting time series that exhibit symmetric random walk characteristics. The weight of the movement term is determined by two parameters: one reflecting the directional accuracy and the other representing the mean absolute increment. The theoretical analysis of the proposed method explains how movement prediction can enhance point forecasting for a random walk-like series. In addition to presenting the relationship between the directional accuracy and the method’s improvement over the original naïve baseline, it also reveals the source of uncertainty in these prediction tasks, which comes from the relationship between the actual and estimated values of the parameters. Furthermore, the simulation results demonstrate that the adjusted naïve forecast can achieve statistically significant improvements over the original naïve forecast, even with directional accuracies slightly greater than 0.50. These findings highlight the significant potential of using movement predictions to address the inherent challenges of point forecasting in a random walk-like time series.
There are several avenues for further research on the foundation of this study. First, while this research focuses on time series with symmetric random walk characteristics, future work could explore applying this movement prediction-based method to time series with nonsymmetric random walk properties, potentially broadening its applicability. Second, it is worth exploring whether the proposed forecasting method can be generalized beyond the one-dimensional case, as many real-world variables exhibit random walk behavior in a high-dimensional space. Moreover, the proposed method uses a fixed increment for simplicity. However, adapting this increment based on the availability of new data may provide more dynamic and accurate forecasts. Investigating how the increments can be adapted and whether such adaptive increments can further refine the forecasting accuracy represents a promising direction for future research.
Acknowledgments
This study did not receive any specific grants from funding agencies in the public, commercial, or not-for-profit sectors.

### v8

@@
 Abstract
-This study introduces a movement-prediction-adjusted naïve forecast, which is the original naïve forecast with the addition of a weighted movement prediction term, in the context of forecasting time series that exhibit symmetric random walk properties.
-The weight of the movement term is determined by two parameters: one reflecting the directional accuracy and the other representing the mean absolute increment.
-The settings of the two parameters involve a trade-off: larger values may yield meaningful gains over the original naïve forecast, whereas smaller values often render the adjusted forecast more reliable.
-This trade-off can be managed by empirically setting the parameters using sliding windows on in-sample data.
-To statistically test the performance of the adjusted naïve forecast under different directional accuracy levels, we used four synthetic time series to simulate multiple forecast scenarios, assuming that for each directional accuracy level, diverse movement predictions were provided.
-The simulation results show that as the directional accuracy increases, the error of the adjusted naïve forecast decreases.
-In particular, the adjusted naïve forecast achieves statistically significant improvements over the original naïve forecast, even under a low directional accuracy of slightly above 0.50.
-This finding implies that the movement-prediction-adjusted naïve forecast can serve as a new optimal point forecast for time series with symmetric random walk characteristics if consistent movement prediction can be provided.
+This study introduces a movement-prediction-adjusted naïve forecast for time series exhibiting symmetric random walk characteristics, which is applicable after accurate movement predictions are available.
+Specifically, the original naïve forecast is adjusted by a weighted movement prediction term, where the weights are determined via two parameters derived from the in-sample data: one based on directional accuracy of the movement prediction and the other on the mean absolute increment of the target series.
+Simulation experiments were conducted across four types of synthetic symmetric random walk series, each with different variance structures.
+For each time series, diverse movement predictions with predefined directional accuracies were randomly generated, and the resulting forecasts were evaluated via the RMSE, MAE, MAPE, and sMAPE metrics.
+The results demonstrated a clear monotonic improvement in the forecast performance as the directional accuracy increased.
+Notably, the adjusted naïve forecast achieved statistically significant improvements even at relatively low directional accuracy levels slightly above 0.50.
+These findings imply that the movement-prediction-adjusted naïve forecast can serve as an effective second-stage method for forecasting symmetric random walk time series when consistent and accurate movement predictions are provided.
 Keywords Naïve forecast  $\cdot$ Movement prediction  $\cdot$ Point forecasting  $\cdot$ Symmetric random walk
@@
 Time series exhibiting random walk characteristics pose significant forecasting challenges (Hamilton,, 2020).
-Therefore, financial time series, known for these characteristics and their practical importance in financial markets, have been extensively studied as test cases for forecasting theories and methods.
+In particular, financial time series, known for these characteristics and their practical importance in financial markets, have been extensively studied as test cases for forecasting theories and methods.
 One prominent theory is the efficient market hypothesis (EMH), which suggests that asset prices incorporate all available information, rendering historical prices unpredictable (Fama,, 1970).
@@
 Moreover, Moosa and Burns, (2014) argued that claims of outperforming the naïve forecast in terms of the RMSE are often misleading, frequently involving introduced dynamics without rigorous statistical validation.
-Despite attempts to use various predictors to break the assumed randomness of financial time series, existing forecasting methods, including statistical, machine learning, and deep learning approaches, typically cannot outperform naïve forecasts (Kilian and Taylor,, 2003; Moosa and Burns,, 2016; Thakkar and Chaudhari,, 2021; Petropoulos et al.,, 2022; Ellwanger and Snudden,, 2023; Hewamalage et al.,, 2023; Zeng et al.,, 2023; Beck et al.,, 2025).
-From the forecaster’s perspective, the unpredictability of random walk-like series could be attributed to incomplete knowledge of the underlying processes controlling the data, ignorance of the various contributing factors, and imperfections in analytical methods (Laplace,, 2012).
-In other words, the randomness of data is not inherent but rather apparent (Crutchfield and Feldman,, 2003).
+Despite numerous attempts to exploit weak structural signals in financial time series via various predictors, existing forecasting methods—including statistical, machine learning, and deep learning approaches—have typically failed to outperform naïve forecasts (Kilian and Taylor,, 2003; Moosa and Burns,, 2016; Thakkar and Chaudhari,, 2021; Petropoulos et al.,, 2022; Ellwanger and Snudden,, 2023; Hewamalage et al.,, 2023; Zeng et al.,, 2023; Beck et al.,, 2025).
+From the forecaster’s perspective, the unpredictability of a variable could be attributed to incomplete knowledge of the underlying processes controlling the data, ignorance of the various contributing factors, and imperfections in analytical methods (Laplace,, 2012).
+In other words, randomness in data is not inherent but rather apparent (Crutchfield and Feldman,, 2003).
 Although useful information may be contained in the input space, it can be overwhelmed by irrelevant data, making it difficult for current regression methods to exploit such information (John et al.,, 1994; Guyon and Elisseeff,, 2003; Verleysen and François,, 2005).
@@
 However, these theories and empirical evidence reveal only the practical difficulty of finding exploitable patterns; they do not eliminate the possibility of discovering any predictability in the data.
-Given the challenge in the point forecasting of random walk-like series, many researchers have shifted their focus from point forecasting to movement prediction, which involves predicting whether the value is increasing or decreasing for the next time step.
+Given the difficulty of generating precise numerical predictions from noisy data in random walk-like time series, researchers have increasingly shifted their focus from point forecasting to movement prediction, which involves predicting whether the value will increase or decrease at the next time step.
 Consequently, the regression task was reduced to a binary classification task.
 For a pure random walk, the probability that future movement will increase or decrease is 0.5, and the mutual information between the predictors and the target is zero (Kraskov et al.,, 2004).
-If the mutual information is greater than zero, it indicates that there is some dependency between the predictors and the target (Cover and Thomas,, 2012).
+For a random walk-like time series using abundant predictors, the mutual information can be greater than zero, indicating that there exists some dependency between the predictors and the target (Cover and Thomas,, 2012).
 Practically, this means that the predictors reduce the uncertainty of the binary outcome.
 Although this does not guarantee high predictive accuracy, it suggests that, theoretically, one can build a model that performs better than random guessing (Ircio et al.,, 2020).
-In contrast, for point forecasting in a random walk-like series, the continuous nature of the target means that the small gain in mutual information only marginally reduces the overall uncertainty and does not sufficiently narrow down the precise numerical result (Beraha et al.,, 2019).
-Movement prediction is often considered easier than point forecasting in financial time series (Taylor,, 2008; Moosa,, 2013).
+In contrast, for point forecasting in a random walk-like series, the continuous nature of the target means that even if a small gain in mutual information exists, it only marginally reduces the overall uncertainty and does not sufficiently narrow down the precise numerical result (Beraha et al.,, 2019).
+Indeed, movement prediction is often considered easier than point forecasting in financial time series (Taylor,, 2008; Moosa,, 2013).
 Studies in the field of behavioral finance suggest that investor sentiment and psychology can lead to certain patterns (De Bondt and Thaler,, 1985; Lee et al.,, 1991; Baker and Wurgler,, 2007; Barberis et al.,, 1998), which could lead to mutual information between the sentiment predictors and the binary target being greater than zero and making future movements predictable.
-Therefore, sentiment indicators are often used to improve the directional accuracy in these tasks.
+Therefore, sentiment indicators are often used to improve the directional accuracy in these movement prediction tasks.
 For example, Weng et al., (2017) incorporated Google news counts and Wikipedia page views alongside historical data to predict the next day’s movement of AAPL stock via support vector machines (SVMs), achieving up to 0.80 directional accuracy.
@@
 When a movement prediction is better than a random guess, it also indicates that the input data contain useful information that has been leveraged effectively to predict future movements.
-Borrowing the "less is more" principle from cognitive science, where reducing information complexity can lead to an improved cognitive process (Gigerenzer and Brighton,, 2009), we can subsequently use the result of the movement prediction to generate the point forecast, using only the binary labels as an additional predictor along with the original naïve forecast, as shown in Figure 1.
-The binary labels can then be regarded as "verified predictors" for the target variable, and no more noise is introduced to affect the final point forecast.
-Unfortunately, the use of movement prediction results only to improve point forecasts remains unexplored in the forecasting community.
-Most studies either stopped at the movement prediction stage or used a direct regression approach to conduct point forecasting.
+Inspired by the ’less is more’ principle, which suggests that reducing informational complexity can improve decision-making (Gigerenzer and Brighton,, 2009), movement prediction can further serve as a simplified predictor in addition to the naïve forecast to generate the final point forecast, as shown in Figure 1.
+This indirect approach requires only the naïve forecast and the movement prediction as predictors, eliminating the need for additional noisy predictors.
+The naïve forecast, which represents the current value, is most closely related to the future value, whereas the movement prediction contributes new directional information not contained within the naïve forecast.
+Unfortunately, the use of movement prediction alone to improve naïve forecasts remains unexplored in the forecasting community.
+Previous studies either perform only movement prediction or apply regression directly to noisy data.
 Although it is often argued that, in many cases, compared with point forecasting, movement prediction is more important for decision-making (Sezer et al.,, 2020), the final decision generally still depends on the estimated value of the target variable.
 For example, in financial markets, the decision could be to trade at a specific price given a predicted future movement.
-Therefore, the decision maker actually conducted the classification-to-regression conversion, albeit implicitly.
-However, how and why movement prediction results can be systematically exploited to improve point forecasting in random walk-like series is largely unknown.
+Thus, even when practitioners rely primarily on predicted movements, they implicitly perform a classification-to-regression conversion by translating the predicted directions into quantitative trading decisions.
+However, there remains a notable lack of systematic approaches for integrating movement prediction outcomes with naïve forecasts in the context of forecasting random walk-like time series.
 Figure 1: Direct and indirect regression approaches
 Under these conditions, this study aimed to explore the potential of using movement prediction to enhance point forecasting in random walk-like time series.
-In particular, the target variable is assumed to follow a symmetric random walk process, which is the basic type of random walk characterized by the absence of drift and equal probabilities of upward and downward movements (Ibe,, 2014).
-By reformulating the symmetric random walk process to incorporate the future movement of the target variable, the point forecast can naturally be expressed as the original naïve forecast adjusted by movement prediction, which is referred to as the movement-prediction-adjusted naïve forecast.
-The weight applied to the movement term is determined by two parameters: one reflecting the directional accuracy and the other representing the mean absolute increment.
-The setting of these two parameters introduces an inherent trade-off and uncertainty into the prediction performance.
-Specifically, larger parameter values suggest potentially meaningful improvements over the original naïve forecast, whereas smaller parameter values indicate a greater likelihood that the adjusted naïve forecast will outperform the original naïve forecast.
-To balance this trade-off, the parameters can be empirically determined by sliding windows on the in-sample data.
+To this end, we propose a second-stage method that effectively integrates movement prediction with the naïve forecast after accurate and consistent movement predictions have been obtained.
+The target variable is assumed to follow a symmetric random walk process, which is the basic type of random walk characterized by the absence of drift and equal probabilities of upward and downward movements (Ibe,, 2014).
+By reformulating the symmetric random walk process to incorporate the future movement of the target variable, the point forecast can naturally be expressed as the original naïve forecast adjusted by the movement prediction term.
+Hence, it is named the movement-prediction-adjusted naïve forecast.
+The weight applied to the movement term is determined by two parameters derived from the in-sample data: one based on the directional accuracy of the movement prediction and the other on the mean absolute increment of the target series.
+The setting of these two parameters involves an inherent trade-off on the prediction performance.
+While larger parameter values indicate a lower likelihood of achieving improvements through adjustment, these improvements can be substantial; smaller values increase the likelihood of improvement, although such improvements may be marginal or negligible compared with the naïve forecast.
+To balance this trade-off, the parameters can be empirically estimated by sliding windows on the in-sample data.
 Since, in practice, movement prediction can exhibit a wide range of directional accuracies, the adjusted naïve forecast should be evaluated across different accuracy levels to examine the impact of directional accuracy on the performance of the proposed method.
@@
 For all the experiments, the evaluation metrics include the RMSE, mean absolute error (MAE), mean absolute percentage error (MAPE), and symmetric mean absolute percentage error (sMAPE), and the original naïve forecast serves as the sole baseline.
-In this experimental setting, it can be determined whether the improvement over the baseline by the proposed method is statistically significant.
-This study makes three significant contributions to the literature.
-First, this study provides new insights into the predictability of random walk-like series, addressing a longstanding challenge in the field.
-When meaningful movement prediction was provided, the point values of the random walk-like series became partially predictable.
-Second, by illuminating the relationship between the actual and estimated parameter values, this study underscores the source of uncertainty in the effectiveness of the proposed method, a critical consideration often underemphasized by purely data-driven approaches.
-Third, the proposed method bridges movement prediction and point forecasting, which are two distinct tasks in time series forecasting.
+In this experimental setting, we can evaluate whether the improvement over the baseline by the proposed method is statistically significant.
+This study makes three significant contributions to the literature:
+  • First, this study offers new insights into the predictability of random walk-like series, presenting a feasible approach to address this longstanding challenge in the field.
+When meaningful movement prediction is available, the point values of these series, which are traditionally viewed as unpredictable, exhibit conditional predictability.
+  • Second, the proposed method bridges movement prediction and point forecasting, which are two distinct tasks in time series forecasting.
 Although the use of binary classification to guide regression is well established, relying solely on the movement prediction of the target variable for its own point forecasting offers a fresh perspective.
+  • Third, this study introduces a clear boundary criterion, which explicitly states that the adjusted naïve forecast outperforms the original naïve forecast when the product of the actual optimal parameter values exceeds half the product of the estimated optimal parameter values.
+This contribution advances forecasting methodologies by clarifying the conditions required for improved performance.
 The remainder of this paper is organized as follows.
-Section 2 introduces the details of the movement-prediction-adjusted naïve forecast, explains how to derive it, and explains why it can outperform the original naïve forecast.
-Section 3 presents a simulation study that assesses the performance of movement-prediction-adjusted naïve forecasts using synthetic datasets in a controlled experimental setting.
+Section 2 introduces the details of the movement-prediction-adjusted naïve forecast, explains how to derive it and under what conditions it can outperform the original naïve forecast.
+Section 3 presents a simulation study that assesses the performance of movement-prediction-adjusted naïve forecasts via synthetic datasets in a controlled experimental setting.
 Section 4 discusses the key findings, implications, and limitations of the study.
 Finally, Section 5 concludes the paper and outlines future research directions.
-2 Movement-Prediction-Adjusted Naïve Forecast
-This section presents the details of the movement-prediction-adjusted naïve forecast for time series that exhibit symmetric random walk characteristics.
-We first derive the general expression of the adjusted naïve forecast based on the definition of symmetric random walks, then explain how to optimize the parameters attached to the movement term, and finally, how to adjust the parameters to improve the forecasting effectiveness and detail the procedure for generating forecasts on the out-of-sample set.
-2.1 General Expression
+2 Method
+This section presents the details of the movement-prediction-adjusted naïve forecast.
+We first derive the general expression of the adjusted naïve forecast based on the definition of symmetric random walks and then explain how to optimize the parameters to improve the forecasting effectiveness and detail the procedure for generating forecasts on the out-of-sample set.
+2.1 General Formulation of the Movement-Prediction-Adjusted Naïve Forecast
 A symmetric random walk is a stochastic process in which the value at time step $t$ is the sum of the value at the previous time step $t-1$ and the value of an independent random increment ${{\epsilon}_{t}}$, which is generally assumed to follow a normal distribution with a mean of 0 and variance of $\sigma_{t}^{2}$ (De Gooijer et al.,, 2017).
@@
 For random walks exhibiting homoscedasticity, the variance $\sigma_{t}^{2}$ remains constant over time, whereas for those exhibiting heteroscedasticity, the variance $\sigma_{t}^{2}$ varies over time.
-The symmetric nature of the random increment ${{\epsilon}_{t}}$ with a mean of 0 ensures that there is no inherent bias in the movement direction of the target variable and that upward and downward movements are equally likely (Ibe,, 2014).
-Furthermore, the random increment $\epsilon_{t}$ can be expressed as a product of two components: a nonnegative random increase $|\epsilon_{t}|$ representing the magnitude of the increment and a sign component ${D}_{t}$, which takes a value of $+1$ for upward movement and $-1$ for downward movement, indicating the direction of the increment between two adjacent points.
+The symmetric nature of the random increment ${{\epsilon}_{t}}$ ensures that there is no inherent bias in the movement direction of the target variable and that upward and downward movements are equally likely (Ibe,, 2014).
+Furthermore, the random increment $\epsilon_{t}$ can be expressed as a product of two components: a nonnegative term $|\epsilon_{t}|$ representing the magnitude of the increment and a sign component ${D}_{t}$, which takes a value of $+1$ for upward movement and $-1$ for downward movement, indicating the direction of the increment between two adjacent points.
 This allows the process to be reformulated as follows:
 ${{y}_{t}}={{y}_{t-1}}+{{D}_{t}}\cdot|\epsilon_{t}|;|\epsilon_{t}|\geq 0$ (2)
-The nonnegative random increment $|\epsilon_{t}|$ can be further expressed as a product of two terms:
+The nonnegative term $|\epsilon_{t}|$ can be further expressed as a product of two terms:
 ${{y}_{t}}={{y}_{t-1}}+{{D}_{t}}\cdot{\theta_{t}}\cdot\bar{\epsilon};{\theta_{t}}\geq 0,\bar{\epsilon}>0,$ (3)
-where $\bar{\epsilon}$ represents the mean absolute increment calculated from the in-sample data, which can be interpreted as a base magnitude of the increments.
+where $\bar{\epsilon}$ represents the mean absolute increment calculated based on the in-sample data, which can be interpreted as a base magnitude of the increments.
 The term $\theta_{t}$ is introduced as a nonnegative parameter that controls the magnitude of the increment at time step $t$.
@@
 (3) implies that the point forecast of the target variable can be generated if ${{D}_{t}}$ can be predicted and both $\theta_{t}$ and $\bar{\epsilon}$ can be estimated.
-Consider a simple scenario in which the magnitude of the increment between two adjacent points is fixed for future time steps; then, $\theta_{t}$ can be set to a constant value.
+Consider a simplified scenario in which the magnitude of the increment between two adjacent points is fixed for future time steps; in this case, $\theta_{t}$ can be set to a constant value.
 Consequently, the point forecast for future time step $t$, denoted as $\hat{y}_{t}$, can be expressed as follows:
@@
 where ${{\hat{D}}_{t}}$ is the movement prediction, $\theta$ is a fixed scalar, and $\theta\in\mathbb{Q}$.
-Therefore, the point forecast generated via Eq.
+Suppose that the movement prediction has been provided, the point forecast generated via Eq.
 (4) can be regarded as an adjustment of the original naïve forecast through a weighted movement prediction term; hence, it is named the movement-prediction-adjusted naïve forecast.
-2.2 Movement Term Weight
+2.2 Parameter Optimization Using In-Sample Data
 According to Eq.
-(4), the weight of the movement term in the adjusted naïve forecast is the product of two parameters, $\theta$ and $\bar{\epsilon}$, both of which can be estimated from the in-sample data.
-Since $\bar{\epsilon}$ is directly computed as the mean absolute increment of the in-sample series, only $\theta$ requires optimization.
-This is typically achieved by minimizing the mean squared error (MSE) of the adjusted naïve forecast for the in-sample set.
-Because the MSE of the original naïve forecast is a fixed value on the in-sample set, minimizing the MSE of the adjusted naïve forecast is mathematically equivalent to maximizing the difference between the MSEs of the original and adjusted naïve forecasts:
+(4), the weight of the movement term in the adjusted naïve forecast is the product of two parameters, $\theta$ and $\bar{\epsilon}$, both of which can be optimized based on the in-sample data.
+Because $\bar{\epsilon}$ is computed directly as the mean absolute increment of the in-sample series, only $\theta$ requires optimization.
+This optimization is typically achieved by minimizing the mean squared error (MSE) of the adjusted naïve forecast within the in-sample set.
+Given that the original naïve forecast’s MSE is constant for the in-sample set, minimizing the adjusted naïve forecast’s MSE is equivalent to maximizing their difference:
 $\min MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\equiv\max\left(MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\right)$ (5)
-Using this approach, we can determine the optimal value of $\theta$ and examine, through the difference between the MSEs, whether there is an improvement over the original naïve forecast under this optimal setting.
-Suppose that the in-sample set has $N$ time steps and that movement prediction $\hat{D}_{t}$ is provided for each time step.
-Within the in-sample set, we assume that the movement prediction yields $n_{\text{in-sample}}^{\text{correct}}$ correct and $n_{\text{in-sample}}^{\text{incorrect}}$ incorrect directions.
-We subsequently used Eq.
-(2) to present the actual value and expand the MSE expression for the original naïve forecast:
-$MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (6)
-$=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t-1}+D_{t}\cdot|\epsilon_{t}|-y_{t-1}\right)^{2}\right)$
-$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\ |\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}\ |\epsilon_{t}|^{2}\right)$
-When using Eq. (2) to represent the actual value and Eq. (4) to represent the predicted value, the MSE of the adjusted naïve forecast on the in-sample set can be expressed as:
-$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$ (7)
-$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(y_{t}-\hat{y}_{t}\right)^{2}+\sum_{t\in\text{incorrect}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
-$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left[(y_{t-1}+D_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{D}_{t}\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}})\right]^{2}\right.$
-$\quad\left.+\sum_{t\in\text{incorrect}}\left[(y_{t-1}+D_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{D}_{t}\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}})\right]^{2}\right)$
-$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
-$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right),$
-where $\theta_{\text{in-sample}}$ represents the parameter $\theta$ estimated from the in-sample set and $\bar{\epsilon}_{\text{in-sample}}$ denotes the mean of $|\epsilon_{t}|$ calculated over the in-sample set.
-Consequently, the difference between the two MSEs on the in-sample set can be derived via Eq.
-(6) and Eq.
-(7), respectively:
-$\Delta\text{MSE}_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (8)
-$=\frac{1}{N}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
-$\quad-\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
-$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
-$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
-$\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
-Based on the law of large numbers (LLN), when both $n_{\text{in-sample}}^{\text{correct}}$ and $n_{\text{in-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
-$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}}$ (9)
-Therefore, Eq. (8) can be simplified as follows:
-$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (10)
-$\approx\frac{1}{N}\left(2\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}})-n_{\text{in-sample}}^{\text{correct}}\cdot\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right.$
-$\quad\left.-2\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}})-n_{\text{in-sample}}^{\text{incorrect}}\cdot\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)$
+Suppose the in-sample set has $N$ time steps, and movement predictions yield $n_{\text{in-sample}}^{\text{correct}}$ correct and $n_{\text{in-sample}}^{\text{incorrect}}$ incorrect directions; then, the directional accuracy of the movement prediction on the in-sample set can be expressed as follows:
+$ACC_{\text{in-sample}}=\frac{n_{\text{in-sample}}^{\text{correct}}}{n_{\text{in-sample}}^{\text{incorrect}}+n_{\text{in-sample}}^{\text{correct}}}$ (6)
+Accordingly, the MSE difference between two types of naïve forecasts is as follows:
+$\Delta MSE_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (7)
 $=(4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}})\cdot\bar{\epsilon}_{\text{in-sample}}^{2},$
-where $ACC_{\text{in-sample}}$ is the directional accuracy of the movement prediction on the in-sample set, which is defined as:
-$ACC_{\text{in-sample}}=\frac{n_{\text{in-sample}}^{\text{correct}}}{n_{\text{in-sample}}^{\text{incorrect}}+n_{\text{in-sample}}^{\text{correct}}}$ (11)
-To maximize the $\Delta MSE_{\text{in-sample}}$, we focus on the expression inside parentheses in the final form of Eq.
-(10), which can be regarded as a function of parameter $\theta_{\text{in-sample}}$:
-$f(\theta_{\text{in-sample}})=4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}}$ (12)
-By differentiating $f(\theta_{\text{in-sample}})$ and setting $f^{\prime}(\theta_{\text{in-sample}})=0$, the critical value of $\theta_{\text{in-sample}}$, denoted as $\theta_{\text{in-sample}}^{*}$, is given by:
+where $\theta_{\text{in-sample}}$ represents the parameter $\theta$ in the adjusted naïve forecast for the in-sample set and $\bar{\epsilon}_{\text{in-sample}}$ denotes the mean of $|\epsilon_{t}|$ calculated over the in-sample set.
+A detailed algebraic derivation of the $\Delta MSE_{\text{in-sample}}$ is provided in Appendix A.
+According to Eq. (7), the optimization reduces to maximizing the following function concerning parameter $\theta_{\text{in-sample}}$:
+$f(\theta_{\text{in-sample}})=4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}},$ (8)
+Differentiating Eq. (8) and setting the derivative to zero yields the optimal solution:
+Proposition 1 (Optimal parameter solution).
+The optimal parameter value $\theta_{\text{in-sample}}^{*}$ that maximizes the MSE difference between the adjusted naïve forecast and the original naïve forecast on the in-sample set is given by:
 $\theta_{\text{in-sample}}^{*}=2\cdot ACC_{\text{in-sample}}-1$ (9)
-Because $f^{\prime\prime}(\theta_{\text{in-sample}})<0$, the function $f(\theta_{\text{in-sample}})$ is concave downward, and the critical point of $f(\theta_{\text{in-sample}})$ is the maximum.
-Therefore, $\theta_{\text{in-sample}}^{*}$ represents the optimal value of parameter $\theta$ in the in-sample set.
-In addition, Eq.
-(13) reveals the relationship between the optimal value of parameter $\theta$ and the directional accuracy of the movement prediction within a fixed sequence length.
-According to Eqs.
-(10) and (13), the improvement over the original naïve forecast on the in-sample set, achieved by the adjusted naïve forecast with optimal parameters, is expressed as follows:
-$\max\Delta MSE_{\text{in-sample}}$ $=\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$ (14)
-According to Eqs.
-(13) and (14), as the directional accuracy increases from 0.5, the value of $f(\theta_{\text{in-sample}})$ increases from 0, and the improvement over the original naïve forecast correspondingly increases from 0.
-Because this improvement is a nonnegative value, the adjusted naïve forecast is at least as good as the original naïve forecast on the in-sample set:
+This result explicitly reveals that the optimal value of the parameter $\theta$ depends linearly on the directional accuracy of the movement predictions for a sequence with a fixed length.
+Since the second derivative of $f(\theta_{\text{in-sample}})$ is negative, this optimal solution is a maximum.
+Substituting Eq. (9) back into Eq. (6) yields:
+Proposition 2 (Performance guarantee).
+Under the optimal parameter setting using $\theta_{\text{in-sample}}^{*}$, the adjusted naïve forecast achieves a nonnegative improvement over the original naïve forecast in terms of the MSE on the in-sample set:
+$\max\Delta MSE_{\text{in-sample}}=\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\geq 0$ (10)
+Proposition 2 confirms that as the directional accuracy increases beyond 0.5, the value of $\theta_{\text{in-sample}}^{*}$ increases from 0, leading to progressively greater improvements over the original naïve forecast.
+At a directional accuracy of 1, the maximum achievable improvement is $\bar{\epsilon}_{\text{in-sample}}^{2}$.
+Thus, with the optimal parameter setting, the adjusted naïve forecast matches or surpasses the performance of the original naïve forecast within the in-sample dataset:
 $MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\leq MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ (11)
-2.3 Forecast on the Out-of-Sample Set
-If we use $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as the estimates of the actual optimal values of $\theta$ and $\bar{\epsilon}$ on the out-of-sample set and suppose that the movement prediction on the out-of-sample set with $M$ time steps yields $m_{\text{out-of-sample}}^{\text{correct}}$ correct directions and $m_{\text{out-of-sample}}^{\text{incorrect}}$ incorrect directions; then, the difference between the two MSEs on the out-of-sample set can be expressed as follows:
-$\Delta\text{MSE}_{\text{out-of-sample}}$ $=\text{MSE}_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (16)
-$=\frac{1}{M}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
-$\quad-\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
-$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
-$=\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
-$\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
-Let $\bar{\epsilon}_{\text{out-of-sample}}$ denote the actual mean absolute increment on the out-of-sample set.
-According to the LLN, when both $m_{\text{out-of-sample}}^{\text{correct}}$ and $m_{\text{out-of-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
-$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}}$ (17)
-Therefore, Eq. (16) can further be expressed as:
+This result underlines the theoretical robustness of the proposed adjustment methodology.
+2.3 Uncertainty Analysis and Practical Generation of Out-of-Sample Forecasts
+We now examine the performance of the adjusted naïve forecast via out-of-sample data.
+Suppose that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are used as the parameters of the adjusted naïve forecast on the out-of-sample set, which contains $M$ time steps.
+Suppose that the movement predictions on this set yield $m_{\text{out-of-sample}}^{\text{correct}}$ correct and $m_{\text{out-of-sample}}^{\text{incorrect}}$ incorrect predictions; then, the directional accuracy of the movement prediction on the out-of-sample set can be expressed as:
+$ACC_{\text{out-of-sample}}=\frac{m_{\text{out-of-sample}}^{\text{correct}}}{m_{\text{out-of-sample}}^{\text{incorrect}}+m_{\text{out-of-sample}}^{\text{correct}}}$ (12)
+Consequently, the MSE difference between two types of naïve forecasts on the out-of-sample is expressed as:
 $\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (13)
-$\approx\frac{1}{M}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}})-m_{\text{out-of-sample}}^{\text{correct}}\cdot\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right.$
-$\quad\left.-2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}})-m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)$
-$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot ACC_{\text{out-of-sample}}\cdot\bar{\epsilon}_{\text{out-of-sample}}-2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}$
-$\quad\cdot(1-ACC_{\text{out-of-sample}})\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
 $=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2},$
-where $ACC_{\text{out-of-sample}}$ is the directional accuracy of the movement prediction on the out-of-sample set, which is defined as:
-$ACC_{\text{out-of-sample}}=\frac{m_{\text{out-of-sample}}^{\text{correct}}}{m_{\text{out-of-sample}}^{\text{incorrect}}+m_{\text{out-of-sample}}^{\text{correct}}},$ (19)
-and $\theta_{\text{out-of-sample}}^{*}$ is the actual optimal value of $\theta$ for the out-of-sample set, which is derived according to Eq. (13):
-$\theta_{\text{out-of-sample}}^{*}=2\cdot ACC_{\text{out-of-sample}}-1$ (20)
-To ensure the effectiveness of the adjusted naïve forecast, the $\Delta MSE_{\text{out-of-sample}}$ should be no less than zero.
+where $\bar{\epsilon}_{\text{out-of-sample}}$ denotes the actual mean absolute increment on the out-of-sample set and $\theta_{\text{out-of-sample}}^{*}$ denotes the actual optimal value of $\theta$ for the out-of-sample set, which is derived according to Proposition 1:
+$\theta_{\text{out-of-sample}}^{*}=2\cdot ACC_{\text{out-of-sample}}-1.$ (14)
+A detailed algebraic expansions of the $\Delta MSE_{\text{out-of-sample}}$ is provided in Appendix B.
+For the adjusted naïve forecast to outperform the original naïve forecast, the $\Delta MSE_{\text{out-of-sample}}$ should be no less than zero.
 Therefore, according to Eq.
-(18), the following inequality must be satisfied:
+(13), the following inequality must hold:
 $\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}\geq\frac{1}{2}\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}$ (15)
-Given that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are specific estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, respectively, inequality (21) can be generalized as:
+Given that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are specific estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, re2spectively, inequality (15) can be generalized as:
+Proposition 3 (Out-of-sample performance criterion).
+The necessary condition for the adjusted naïve forecast to achieve an improvement over the original naïve forecast in terms of the MSE on the out-of-sample set is as follows:
 $\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}\geq\frac{1}{2}\cdot\hat{\theta}_{\text{out-of-sample}}\cdot\hat{\bar{\epsilon}}_{\text{out-of-sample}},$ (16)
 where $\hat{\theta}_{\text{out-of-sample}}$ represents the estimate of $\theta_{\text{out-of-sample}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ represents the estimate of $\bar{\epsilon}_{\text{out-of-sample}}$.
-Since the values of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ cannot be determined in advance, whether the inequality (22) can be satisfied is unknown, even when the optimal parameters derived from the in-sample set are used.
-Therefore, inequality (22) reveals the source of the uncertainty in the effectiveness of the adjusted naïve forecast, which depends on whether the product of the actual parameters exceeds half the product of the estimated parameters.
-We can regard $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ as two variables, each of which follows a distribution with a nonnegative range.
-Inequality (22) represents an event whose probability of occurrence depends on the estimated values of the two variables.
-When both estimates are set to small values near zero, the probability that inequality (22) is satisfied approaches one.
-However, in such a setting, the improvement over the original naïve forecast by the adjusted naïve forecast on the out-of-sample set also diminishes.
-Therefore, inequality (22) also reflects a trade-off between maximizing the probability that the adjusted naïve forecast outperforms the original forecast and achieving a meaningful level of improvement.
-In addition to using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, which may result in overfitting, a more conservative approach can be used to set the estimated values of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$:
+Proposition 3 establishes a clear boundary criterion when an adjusted naïve forecast can outperform the original naïve forecast, specifically depending on whether the product of the actual optimal parameter values exceeds half the product of the estimated optimal parameter values.
+We can treat $\theta_{\text{out-of-sample}}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ as random variables, each following a distribution defined over nonnegative values.
+Accordingly, inequality (16) represents a probabilistic event, the likelihood of which depends on the realized values of these variables.
+Since $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ cannot be known in advance, the fulfillment of inequality (16) is uncertain.
+Larger parameter values have a lower likelihood of yielding improvements, although such improvements can be substantial, whereas smaller parameter values increase the likelihood of improvement, albeit potentially marginal or negligible, compared with the naïve forecast.
+Therefore, this boundary criterion highlights a critical trade-off between the likelihood of achieving improvements and the potential magnitude of these improvements.
+In addition to using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, a more conservative approach can be used to generate the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$:
   1. 1.
@@
 (9).
-As the window moves, a collection of $\theta$ values is generated, denoted as $\Theta_{\text{set}}$, whose bounds can be regarded as the estimated range of $\theta_{\text{out-of-sample}}^{*}$.
+As the window moves, a collection of $\theta$ values is generated, denoted as $\Theta_{\text{set}}$, whose bounds can be regarded as the empirical range of $\theta_{\text{out-of-sample}}^{*}$.
   3. 3.
-    Similarly, a sliding window of the same length is applied to calculate a set of $\bar{\epsilon}$ values from the in-sample set, denoted as $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, providing an estimated value range of $\bar{\epsilon}_{\text{out-of-sample}}$.
+    Similarly, a sliding window of the same length is applied to calculate a set of $\bar{\epsilon}$ values from the in-sample set, denoted as $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, providing an empirical value range of $\bar{\epsilon}_{\text{out-of-sample}}$.
   4. 4.
@@
 end for
-6. Set the initial forecast:
+6. Forecast the out-of-sample set:
 $Y_{\text{last}}\leftarrow Y_{\text{in-sample}}[\text{end}]$ //    Start from the last element of the in-sample set
-7. Forecast the out-of-sample set:
 for $i\leftarrow 1$ to $\text{length}(Y_{\text{out-of-sample}})-1$ do
@@
 3 Simulation
-This section presents a simulation study using four synthetic time series to evaluate whether the adjusted naïve forecast outperforms the original naïve forecast under different directional accuracy levels; therefore, movement prediction is assumed to have been provided.
-To statistically test the performance of the proposed method on an out-of-sample set, multiple movement predictions were randomly generated for each directional accuracy level.
+This section presents a simulation study using four synthetic time series to evaluate whether the adjusted naïve forecast outperforms the original naïve forecast when movement predictions with different directional accuracy levels are provided.
+The performance of the proposed method under each directional accuracy level was evaluated based on the out-of-sample set of each time series.
 3.1 Datasets and Setup
 The dataset includes four synthetic time series, each of which represents a symmetric random walk with a certain type of variance, as summarized in Table 1.
-Each time series, consisting of 2,500 steps, used a fixed random seed (seed = 1) in the Python environment, as shown in Figure 2.
+Each time series, consisting of 2,500 steps, was generated using a fixed random seed (seed = 1) in the Python environment, as shown in Figure 2.
 To maintain all values within the positive range, an initial positive offset of 10000 units was applied to each series.
@@
 In addition, Figure 3 shows the first-order differences of each time series to visually demonstrate the properties of the increments, and Figure 4 shows the histograms of the first-order differences for all time series, highlighting the symmetric nature of the time series.
-Table 1: Synthetic time series summary
+Table 1: Variance of synthetic random walks
 Variance Type Expression Parameter Range
@@
 Figure 4: Histograms of first-order differences of four synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
-For each time series, the adjusted naïve forecast on the out-of-sample set was generated via Algorithm 1, which uses the values of $\theta$ determined by different levels of predefined directional accuracies.
-The performance of each adjusted naïve forecast result was evaluated using four metrics: the RMSE, MAE, MAPE, and sMAPE.
+For each time series, the adjusted naïve forecast on the out-of-sample set was generated via Algorithm 1.
+The performance of each adjusted naïve forecast result was evaluated via four metrics: the RMSE, MAE, MAPE, and sMAPE.
 The RMSE, derived from the square root of the MSE, emphasizes larger errors, which makes it suitable for identifying significant deviations in forecasts.
@@
 To statistically test the performance of the adjusted naïve forecast, for each synthetic time series at a specified directional accuracy level, the process to generate the adjusted naïve forecast was repeated 100 times.
-Each repetition used a unique movement prediction sequence at the same directional accuracy level, which was generated by randomly replacing a subset of actual movements in the out-of-sample set with incorrect movements.
+Each repetition uses a unique movement prediction sequence at the same directional accuracy level, which is generated by randomly replacing a subset of actual movements in the out-of-sample set with incorrect movements.
 The directional accuracy level varies in two ways: initially, it changes from 0.50 to 0.56 in increments of 0.01, and then it changes from 0.55 to 1.00 in larger increments of 0.05.
@@
 3.2 Results and Analysis
-Because the directional accuracy for the out-of-sample set is predefined in each simulation process, only the value of $\bar{\epsilon}_{\text{out-of-sample}}$ needs to be estimated.
-Figure 5 illustrates the distribution of the values in the $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, generated on the basis of sliding windows across the first-order differences of the in-sample set.
-The vertical reference line in Figure 5 represents the actual $\bar{\epsilon}_{\text{out-of-sample}}$ value of the out-of-sample set.
+Because the directional accuracy for the out-of-sample set is predefined in each simulation process, the value of $\theta_{\text{out-of-sample}}^{*}$ can be derived directly via Eq.
+(14), and only the value of $\bar{\epsilon}_{\text{out-of-sample}}$ needs to be estimated.
+Figure 5 illustrates the distribution of the values in the $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, generated based on the sliding windows across the first-order differences of the in-sample set.
+The vertical reference line in Figure 5 represents the actual value of $\bar{\epsilon}_{\text{out-of-sample}}$.
 In the four cases, the minimum value of $\boldsymbol{\bar{\epsilon}}_{\text{set}}$ is generally an underestimate of $\bar{\epsilon}_{\text{out-of-sample}}$, except for the second case, where the synthetic time series is a random walk with a decreasing trend of linear variance.
-However, the estimated value remains close to the true value.
-Under the assumption that directional accuracy is consistent across the in-sample and out-of-sample sets, this estimation ensures that inequality (22) is satisfied.
+In general, the estimated value remains close to the true value.
+Under the assumption that the directional accuracy is consistent across the in-sample and out-of-sample sets, this estimation ensures that inequality (16) is satisfied.
 Figure 5: Comparison of the true $\bar{\epsilon}_{\text{out-of-sample}}$ with $\boldsymbol{\bar{\epsilon}}_{\text{set}}$: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
 Furthermore, comparisons of the two types of naïve forecasts based on the four evaluation metrics at low accuracy levels are illustrated in Figures 6, 7, 8, and 9.
-Each box plot in each figure represents the distribution of 100 independent metric values at a specific directional accuracy level, with the metric value of the original naïve forecast shown as a reference line.
+Each box plot in each figure represents the distribution of 100 independent values for one metric at a specific directional accuracy level, with the metric value of the original naïve forecast shown as a reference line.
 According to Figure 6, at low accuracy levels, such as 0.51, the RMSE of the adjusted naïve forecast is comparable to or slightly better than that of the original naïve forecast, as shown by the proximity of the median of each box plot to the reference line.
@@
 As summarized in Table 2, at a directional accuracy level of 0.51, the adjusted naïve forecast shows mixed results with respect to statistical significance.
-Although significant improvements were observed over the original naïve forecast for the Type 4 series (randomly varying variance, p = 0.002), the results for the Type 1 (constant variance, p = 0.915), Type 2 (linear variance trend, p = 0.781), and Type 3 (cyclic variance, p = 0.812) series were not statistically significant, indicating that an accuracy level of 0.51 may not consistently produce meaningful improvements across all types of time series.
+Although significant improvements were observed over the original naïve forecast for the Type 4 series (randomly varying variance, p = 0.002), the results for the Type 1 (constant variance, p = 0.915), Type 2 (linear variance trend, p = 0.781), and Type 3 (cyclic variance, p = 0.812) series were not statistically significant, indicating that the adjusted naïve forecast under an accuracy level of 0.51 may not consistently produce meaningful improvements across all types of time series.
 As the accuracy increased beyond 0.51, significant reductions in the RMSE (p < 0.001) were observed for all types of time series, thereby demonstrating the reliability and effectiveness of the adjusted naïve forecast at higher accuracy levels.
@@
 According to Figures 10, 11, 12, and 13, at higher accuracy levels ranging from 0.55–1.00, the adjusted naïve forecast consistently outperforms the original naïve forecast in terms of the four metrics.
-In particular, the improvement over the original naïve forecast becomes more pronounced as the directional accuracy improves.
-This finding demonstrates that highly accurate movement predictions can be effectively translated into improved point forecasts in the context of forecasting random walk-like time series.
-However, the improvement over the original naïve forecast is limited, because the proposed method allows a random walk-like series to be predicted to a certain extent, but the predicted value cannot be infinitely close to the actual value.
+In particular, the improvement over the original naïve forecast becomes substantial as the directional accuracy improves.
+This finding demonstrates that the proposed method can effectively translate highly accurate movement predictions into improved point forecasts in the context of forecasting random walk-like time series.
+However, the improvement over the original naïve forecast has a limit.
+The proposed method allows a random walk-like series to be predicted to a certain extent, but the predicted value cannot be infinitely close to the actual value.
 Figure 10: RMSEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
@@
 In general, when the directional accuracy was 0.5, the adjusted naïve forecast did not provide an improvement over the original naïve forecast.
-However, as the directional accuracy increases, the performance of the adjusted naïve forecast improves.
+As the directional accuracy increases, the performance of the adjusted naïve forecast improves.
+The simulation results also demonstrate that the effectiveness of the proposed method is unaffected by the variance type of the synthetic time series, as evidenced by the consistent performance of the adjusted naïve forecast across all synthetic series.
 This result is consistent with the theoretical analysis of the proposed method, which highlights its robustness and potential applicability to a wide range of time series that exhibit symmetric random walk characteristics, including those with homoscedasticity as well as high-volatility scenarios.
@@
 This section discusses the implications of the findings and potential limitations of this study.
-First, this study demonstrates an indirect regression approach that uses only movement predictions and the original naïve forecast for a random walk-like time series.
-The theoretical analysis and simulation results indicate that as the directional accuracy increases, the improvement over the original naïve forecast also increases, suggesting that improving movement predictions may be more critical than developing direct regression models in this context.
-It also highlights a fundamental trade-off between maximizing the probability that the method outperforms the original naïve forecast and achieving a meaningful level of improvement, reflecting an inherent uncertainty in prediction tasks that purely data-driven approaches may sometimes overlook.
-Second, many real-world forecasting scenarios require the integration of predictions from different data types to support decision-making (Bishop and Nasrabadi,, 2006; Timmermann,, 2006).
-However, it is difficult to combine binary and continuous forecasts.
-The proposed method can be regarded as a feasible approach to convert binary movement predictions into continuous point forecasts.
-This conversion facilitates the comprehensive use of diverse predictions and can enhance decision-making in practice.
-Third, this study focuses on the theoretical predictability of random walk-like time series rather than exploiting this predictability for profit.
-Although certain structural properties or external signals may slightly improve movement predictions and thereby point forecasts, the forecasting result does not necessarily translate into economic gains, especially when transaction costs, market frictions, and slippage are considered.
-Finally, because the aim of this study was to show how incorporating movement predictions can enhance naïve forecasts rather than how to generate movement predictions, we relied on simulations rather than real-world examples.
-Notably, the effectiveness of the adjusted naïve forecast depends on the consistent directional accuracy across different segments of the time series.
-Movement prediction models prone to overfitting may show a decline in directional accuracy on the out-of-sample set, limiting the actual improvement over the naïve forecast using the proposed method.
+First, this study demonstrates an indirect regression approach that leverages movement predictions in conjunction with naïve forecasts for time series exhibiting random walk-like behavior.
+The theoretical analysis and simulation results suggest that accurate movement prediction can be systematically translated into improved point forecasts.
+This highlights the potential value of incorporating directional information into simple baseline models and suggests that enhancing movement prediction may, in some contexts, be more effective than developing increasingly complex direct regression models.
+Second, in many real-world forecasting scenarios, decision-making often involves integrating predictions with heterogeneous data types, such as combining binary classifications and continuous regressions.
+However, such integration can be technically challenging.
+The proposed method provides a practical means of converting binary movement predictions into continuous point forecasts for random walk-like time series, thereby enabling more informed decision-making in applied forecasting contexts.
+Third, the goal of this study is to examine the theoretical possibility of improving forecasts for random walk-like time series by incorporating directional signals rather than disputing the broader validity of the naïve forecast as the optimal benchmark.
+While the simulation implies that even weak directional signals can be exploited to enhance point forecasts, these improvements can be modest and context dependent.
+The adjusted naïve forecast serves as a second-stage method only after movement prediction is provided.
+Moreover, such improvements do not necessarily translate into economic gains, particularly when real-world constraints such as transaction costs, slippage, and market frictions are considered.
+Finally, because the focus of this study is on the theoretical impact of directional information rather than on modeling movement predictions themselves, we rely on controlled simulations rather than empirical data.
+The effectiveness of the proposed adjustment depends on the consistency of the directional accuracy across different segments of the series.
+In practice, movement prediction models susceptible to overfitting may suffer from deteriorated performance on out-of-sample data, thereby limiting the actual benefit of incorporating such predictions into the original naïve forecast.
 5 Conclusion
 This study introduces a movement-prediction-adjusted naïve forecast, which is a linear combination of movement prediction and the original naïve forecast, in the challenging context of forecasting time series that exhibit symmetric random walk characteristics.
-The weight of the movement term is determined by two parameters: one reflecting the directional accuracy and the other representing the mean absolute increment.
-The theoretical analysis of the proposed method explains how movement prediction can enhance point forecasting for a random walk-like series.
-In addition to presenting the relationship between the directional accuracy and the method’s improvement over the original naïve baseline, it also reveals the source of uncertainty in these prediction tasks, which comes from the relationship between the actual and estimated values of the parameters.
+The weight of the movement term is determined by two parameters derived from the in-sample data: one based on the directional accuracy of the movement prediction and the other on the mean absolute increment of the target series.
+The theoretical analysis of the proposed method explains how accurate movement prediction can be exploited to enhance point forecasting for a random walk-like series.
+It also defines a boundary criterion stating that the adjusted naïve forecast outperforms the original naïve forecast when the product of the actual optimal parameter values exceeds half the product of their estimated values.
 Furthermore, the simulation results demonstrate that the adjusted naïve forecast can achieve statistically significant improvements over the original naïve forecast, even with directional accuracies slightly greater than 0.50.
@@
 This study did not receive any specific grants from funding agencies in the public, commercial, or not-for-profit sectors.
+Appendix A. Algebraic Derivation of $\Delta MSE_{\text{in-sample}}$
+According to Eq. (2), which presents the actual value of the target variable, the MSE expression for the original naïve forecast can be expressed as:
+$MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (A1)
+$=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t-1}+D_{t}\cdot|\epsilon_{t}|-y_{t-1}\right)^{2}\right)$
+$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\ |\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}\ |\epsilon_{t}|^{2}\right)$
+Similarly, by substituting Eq. (2) for actual values and Eq. (4) for predicted values, the MSE of the adjusted naïve forecast can be expressed as follows:
+$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$ (A2)
+$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(y_{t}-\hat{y}_{t}\right)^{2}+\sum_{t\in\text{incorrect}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
+$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left[(y_{t-1}+D_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{D}_{t}\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}})\right]^{2}\right.$
+$\quad\left.+\sum_{t\in\text{incorrect}}\left[(y_{t-1}+D_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{D}_{t}\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}})\right]^{2}\right)$
+$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
+$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right),$
+Hence, the MSE difference between the adjusted naïve forecast and the original naïve forecast on the in-sample set can be derived via Eq. (A1) and Eq. (A2):
+$\Delta\text{MSE}_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (A3)
+$=\frac{1}{N}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
+$\quad-\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
+$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
+$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
+$\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
+Based on the law of large numbers (LLN), when both $n_{\text{in-sample}}^{\text{correct}}$ and $n_{\text{in-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
+$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}}$ (A4)
+Therefore, Eq. (A3) can be simplified as follows::
+$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (A5)
+$\approx\frac{1}{N}\left(2\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}})-n_{\text{in-sample}}^{\text{correct}}\cdot\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right.$
+$\quad\left.-2\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}})-n_{\text{in-sample}}^{\text{incorrect}}\cdot\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)$
+$=(4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}})\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
+Appendix B. Algebraic Derivation of $\Delta MSE_{\text{out-of-sample}}$
+When using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as parameters for the adjusted naïve forecast on the out-of-sample set, the difference in MSE between the adjusted and original naïve forecasts on the out-of-sample set can be expressed as follows:
+$\Delta\text{MSE}_{\text{out-of-sample}}$ $=\text{MSE}_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (B1)
+$=\frac{1}{M}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
+$\quad-\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
+$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
+$=\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
+$\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
+According to the LLN, when both $m_{\text{out-of-sample}}^{\text{correct}}$ and $m_{\text{out-of-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
+$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}}$ (B2)
+Therefore, Eq. (B1) can further be expressed as:
+$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (B3)
+$\approx\frac{1}{M}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}})-m_{\text{out-of-sample}}^{\text{correct}}\cdot\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right.$
+$\quad\left.-2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}})-m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)$
+$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot ACC_{\text{out-of-sample}}\cdot\bar{\epsilon}_{\text{out-of-sample}}-2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}$
+$\quad\cdot(1-ACC_{\text{out-of-sample}})\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
+$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$

### v9

@@
-Movement-Prediction-Adjusted Naïve Forecast
+Movement-Prediction-Adjusted Naive Forecast
 Abstract
-This study introduces a movement-prediction-adjusted naïve forecast for time series exhibiting symmetric random walk characteristics, which is applicable after accurate movement predictions are available.
-Specifically, the original naïve forecast is adjusted by a weighted movement prediction term, where the weights are determined via two parameters derived from the in-sample data: one based on directional accuracy of the movement prediction and the other on the mean absolute increment of the target series.
-Simulation experiments were conducted across four types of synthetic symmetric random walk series, each with different variance structures.
-For each time series, diverse movement predictions with predefined directional accuracies were randomly generated, and the resulting forecasts were evaluated via the RMSE, MAE, MAPE, and sMAPE metrics.
-The results demonstrated a clear monotonic improvement in the forecast performance as the directional accuracy increased.
-Notably, the adjusted naïve forecast achieved statistically significant improvements even at relatively low directional accuracy levels slightly above 0.50.
-These findings imply that the movement-prediction-adjusted naïve forecast can serve as an effective second-stage method for forecasting symmetric random walk time series when consistent and accurate movement predictions are provided.
-Keywords Naïve forecast  $\cdot$ Movement prediction  $\cdot$ Point forecasting  $\cdot$ Symmetric random walk
+In financial time series forecasting, surpassing the naive forecast is challenging due to the randomness in the data.
+To address this challenge, this study proposes a novel forecast combination method, the movement-prediction-adjusted naive forecast (MPANF), which is designed to improve point forecasts beyond the naive baseline.
+Specifically, MPANF integrates two forecasting components: a naive forecast and a movement prediction.
+The final forecast is generated by adjusting the naive forecast with a movement prediction term, the weight of which is the product of two in-sample quantities: one is a coefficient determined from the movement prediction accuracy and the other is the mean absolute increment.
+The performance of MPANF was evaluated on eight financial time series via standard metrics, including the RMSE, MAE, MAPE, and sMAPE.
+Under modest movement prediction accuracy slightly above 0.55, MPANF generally outperforms common benchmarks such as the naive forecast, naive forecast with drift, integrated moving average of order (1,1) (IMA(1,1)), and linear regression.
+These findings suggest that MPANF can serve as an effective second-stage method when reliable movement predictions are available.
+Keywords Financial time series  $\cdot$ Point forecasting  $\cdot$ Movement prediction  $\cdot$ Naive forecast  $\cdot$ Second-stage forecast combination
 1 Introduction
-Time series forecasting plays a crucial role in many decision-making processes.
-The widespread demand for accurate point forecasts across various domains has attracted considerable attention from researchers and practitioners.
-However, not all time series are equally predictable.
-Time series exhibiting random walk characteristics pose significant forecasting challenges (Hamilton,, 2020).
-In particular, financial time series, known for these characteristics and their practical importance in financial markets, have been extensively studied as test cases for forecasting theories and methods.
-One prominent theory is the efficient market hypothesis (EMH), which suggests that asset prices incorporate all available information, rendering historical prices unpredictable (Fama,, 1970).
-For a random walk-like series, the naïve forecast, which assumes that the future value at the next time step is equal to the current value, is often regarded as the optimal point forecast.
-In support of the EMH, Moosa, (2013) reported that as the volatility of financial time series increases, the root mean square error (RMSE) of the forecasting models increases more quickly than that of the naïve forecast.
-Moreover, Moosa and Burns, (2014) argued that claims of outperforming the naïve forecast in terms of the RMSE are often misleading, frequently involving introduced dynamics without rigorous statistical validation.
-Despite numerous attempts to exploit weak structural signals in financial time series via various predictors, existing forecasting methods—including statistical, machine learning, and deep learning approaches—have typically failed to outperform naïve forecasts (Kilian and Taylor,, 2003; Moosa and Burns,, 2016; Thakkar and Chaudhari,, 2021; Petropoulos et al.,, 2022; Ellwanger and Snudden,, 2023; Hewamalage et al.,, 2023; Zeng et al.,, 2023; Beck et al.,, 2025).
+Financial time series forecasting plays a crucial role in many decision-making processes.
+The widespread demand for accurate point forecasts in the financial market has attracted considerable attention from researchers and practitioners.
+However, although financial time series have been extensively studied as test cases for forecasting theories and methods, randomness in the data often poses significant challenges in generating accurate forecasts (Hamilton,, 2020).
+One prominent theory on this challenge is the efficient market hypothesis (EMH), which suggests that asset prices incorporate all available information, rendering historical prices unpredictable (Fama,, 1970).
+The naive forecast, or persistent forecast, which assumes that the next value is equal to the current value, is frequently regarded as the optimal benchmark for these tasks.
+In support of the EMH, Moosa, (2013) reported that as the volatility of financial time series increases, the root mean square error (RMSE) of the forecasting models increases more quickly than that of the naive forecast.
+Moreover, Moosa and Burns, (2014) argued that claims of outperforming the naive forecast in terms of the RMSE are often misleading, frequently involving introduced dynamics without rigorous statistical validation.
+Despite numerous attempts to exploit weak structural signals in financial time series, existing forecasting methods—including statistical, machine learning, and deep learning approaches—have typically failed to outperform naive forecasts (Kilian and Taylor,, 2003; Moosa and Burns,, 2016; Thakkar and Chaudhari,, 2021; Petropoulos et al.,, 2022; Ellwanger and Snudden,, 2023; Hewamalage et al.,, 2023; Zeng et al.,, 2023; Beck et al.,, 2025).
 From the forecaster’s perspective, the unpredictability of a variable could be attributed to incomplete knowledge of the underlying processes controlling the data, ignorance of the various contributing factors, and imperfections in analytical methods (Laplace,, 2012).
 In other words, randomness in data is not inherent but rather apparent (Crutchfield and Feldman,, 2003).
-Although useful information may be contained in the input space, it can be overwhelmed by irrelevant data, making it difficult for current regression methods to exploit such information (John et al.,, 1994; Guyon and Elisseeff,, 2003; Verleysen and François,, 2005).
+Although useful information may be present in the input space, it can be overwhelmed by irrelevant information, making it difficult for the regression method to exploit the useful content effectively (John et al.,, 1994; Guyon and Elisseeff,, 2003; Verleysen and François,, 2005).
 Moreover, identifying the predictors on which the target variable truly depends is a complex task (Granger,, 1969; Stock and Watson,, 2002).
-However, these theories and empirical evidence reveal only the practical difficulty of finding exploitable patterns; they do not eliminate the possibility of discovering any predictability in the data.
-Given the difficulty of generating precise numerical predictions from noisy data in random walk-like time series, researchers have increasingly shifted their focus from point forecasting to movement prediction, which involves predicting whether the value will increase or decrease at the next time step.
-Consequently, the regression task was reduced to a binary classification task.
-For a pure random walk, the probability that future movement will increase or decrease is 0.5, and the mutual information between the predictors and the target is zero (Kraskov et al.,, 2004).
-For a random walk-like time series using abundant predictors, the mutual information can be greater than zero, indicating that there exists some dependency between the predictors and the target (Cover and Thomas,, 2012).
-Practically, this means that the predictors reduce the uncertainty of the binary outcome.
+Nevertheless, these theories and empirical evidence reveal only the practical difficulty of finding exploitable patterns; they do not eliminate the possibility of discovering any predictability in the data.
+In a forecasting task, if the mutual information between the predictors and the target variable is greater than zero, it indicates that a dependency between them exists (Kraskov et al.,, 2004; Cover and Thomas,, 2012).
+In practical terms, this means that the predictors can reduce uncertainty in binary forecasting tasks, such as determining whether the target value will rise or fall in future time steps.
 Although this does not guarantee high predictive accuracy, it suggests that, theoretically, one can build a model that performs better than random guessing (Ircio et al.,, 2020).
-In contrast, for point forecasting in a random walk-like series, the continuous nature of the target means that even if a small gain in mutual information exists, it only marginally reduces the overall uncertainty and does not sufficiently narrow down the precise numerical result (Beraha et al.,, 2019).
+In contrast, for a point forecasting task, the continuous nature of the target means that even if a small gain in mutual information exists, it only marginally reduces the overall uncertainty and does not sufficiently narrow down the precise numerical output (Beraha et al.,, 2019).
+Given the difficulty of generating precise numerical predictions from noisy data in financial time series, researchers have increasingly shifted their focus from point forecasting to movement prediction.
 Indeed, movement prediction is often considered easier than point forecasting in financial time series (Taylor,, 2008; Moosa,, 2013).
-Studies in the field of behavioral finance suggest that investor sentiment and psychology can lead to certain patterns (De Bondt and Thaler,, 1985; Lee et al.,, 1991; Baker and Wurgler,, 2007; Barberis et al.,, 1998), which could lead to mutual information between the sentiment predictors and the binary target being greater than zero and making future movements predictable.
-Therefore, sentiment indicators are often used to improve the directional accuracy in these movement prediction tasks.
+Behavioral finance studies further suggest that investor sentiment and psychological factors can generate predictable movement patterns (De Bondt and Thaler,, 1985; Lee et al.,, 1991; Baker and Wurgler,, 2007; Barberis et al.,, 1998).
+Consequently, predictors such as sentiment indicators are often used to improve the directional accuracy in these movement prediction tasks.
 For example, Weng et al., (2017) incorporated Google news counts and Wikipedia page views alongside historical data to predict the next day’s movement of AAPL stock via support vector machines (SVMs), achieving up to 0.80 directional accuracy.
-Ma et al., (2023) proposed a multisource aggregated classification (MAC) method for the prediction of stock price movement, incorporating the numerical features and market-driven news sentiments of target stocks, as well as the news sentiments of their related stocks, and the directional accuracy can reach as high as 0.70.
-Wang et al., (2025) used multimodal data that included news data to train a hybrid model consisting of a long short memory network (LSTM) and transformer for the prediction of stock movement, and the directional accuracy fell within the range of 0.57–0.87.
-Depending on the complexity of the markets and availability of data, the directional accuracy of financial time series typically ranges between 0.55 and 0.80 (Bustos and Pomares-Quimbaya,, 2020).
-When a movement prediction is better than a random guess, it also indicates that the input data contain useful information that has been leveraged effectively to predict future movements.
-Inspired by the ’less is more’ principle, which suggests that reducing informational complexity can improve decision-making (Gigerenzer and Brighton,, 2009), movement prediction can further serve as a simplified predictor in addition to the naïve forecast to generate the final point forecast, as shown in Figure 1.
-This indirect approach requires only the naïve forecast and the movement prediction as predictors, eliminating the need for additional noisy predictors.
-The naïve forecast, which represents the current value, is most closely related to the future value, whereas the movement prediction contributes new directional information not contained within the naïve forecast.
-Unfortunately, the use of movement prediction alone to improve naïve forecasts remains unexplored in the forecasting community.
-Previous studies either perform only movement prediction or apply regression directly to noisy data.
-Although it is often argued that, in many cases, compared with point forecasting, movement prediction is more important for decision-making (Sezer et al.,, 2020), the final decision generally still depends on the estimated value of the target variable.
-For example, in financial markets, the decision could be to trade at a specific price given a predicted future movement.
-Thus, even when practitioners rely primarily on predicted movements, they implicitly perform a classification-to-regression conversion by translating the predicted directions into quantitative trading decisions.
-However, there remains a notable lack of systematic approaches for integrating movement prediction outcomes with naïve forecasts in the context of forecasting random walk-like time series.
-Figure 1: Direct and indirect regression approaches
-Under these conditions, this study aimed to explore the potential of using movement prediction to enhance point forecasting in random walk-like time series.
-To this end, we propose a second-stage method that effectively integrates movement prediction with the naïve forecast after accurate and consistent movement predictions have been obtained.
-The target variable is assumed to follow a symmetric random walk process, which is the basic type of random walk characterized by the absence of drift and equal probabilities of upward and downward movements (Ibe,, 2014).
-By reformulating the symmetric random walk process to incorporate the future movement of the target variable, the point forecast can naturally be expressed as the original naïve forecast adjusted by the movement prediction term.
-Hence, it is named the movement-prediction-adjusted naïve forecast.
-The weight applied to the movement term is determined by two parameters derived from the in-sample data: one based on the directional accuracy of the movement prediction and the other on the mean absolute increment of the target series.
-The setting of these two parameters involves an inherent trade-off on the prediction performance.
-While larger parameter values indicate a lower likelihood of achieving improvements through adjustment, these improvements can be substantial; smaller values increase the likelihood of improvement, although such improvements may be marginal or negligible compared with the naïve forecast.
-To balance this trade-off, the parameters can be empirically estimated by sliding windows on the in-sample data.
-Since, in practice, movement prediction can exhibit a wide range of directional accuracies, the adjusted naïve forecast should be evaluated across different accuracy levels to examine the impact of directional accuracy on the performance of the proposed method.
-Therefore, we use synthetic time series to simulate multiple forecasting scenarios, assuming that for each directional accuracy level, diverse movement predictions are provided.
-For all the experiments, the evaluation metrics include the RMSE, mean absolute error (MAE), mean absolute percentage error (MAPE), and symmetric mean absolute percentage error (sMAPE), and the original naïve forecast serves as the sole baseline.
-In this experimental setting, we can evaluate whether the improvement over the baseline by the proposed method is statistically significant.
+Ma et al., (2023) proposed a multisource aggregated classification (MAC) method for the prediction of stock price movement, incorporating the numerical features and market-driven news sentiments of target stocks, and the directional accuracy reached 0.70.
+Wang et al., (2025) used multimodal data that included news data to train a hybrid model consisting of a long short memory network (LSTM) and a transformer for the prediction of stock movement, and the directional accuracy fell within the range of 0.57–0.87.
+Depending on the complexity of the markets and availability of data, the movement prediction accuracy of financial time series typically ranges between 0.55 and 0.80 (Bustos and Pomares-Quimbaya,, 2020).
+These accuracy levels, which are consistently above the 0.50 threshold of random guessing, confirm that financial time series movement prediction is feasible.
+This naturally raises the question of whether the directional information captured by movement prediction can be systematically incorporated into point forecasting to increase accuracy.
+A forecasting framework for this purpose should align with the “less is more” principle, which posits that reducing informational complexity can enhance decision-making (Gigerenzer and Brighton,, 2009).
+Accordingly, it should rely solely on the naive forecast and the movement prediction as inputs, thereby avoiding additional noisy predictors (see Figure 1).
+The naive forecast, which represents the current value, is most closely related to the future value, whereas the movement prediction contributes new directional information not contained within the naive forecast.
+Unfortunately, previous studies either focused exclusively on movement prediction or applied regression directly to noisy data.
+Although forecast combination and consensus forecasting are well established, they typically blend multiple point forecasts rather than use directional signals to improve point forecast accuracy.
+The use of movement prediction alone for this purpose remains unexplored, leaving a notable gap in systematic approaches for integrating movement prediction outcomes with naive forecasts in time series forecasting.
+Figure 1: Direct and indirect frameworks for point forecasting.
+To investigate the use of movement prediction for improving point forecasting in financial time series, we propose the movement-prediction-adjusted naive forecast (MPANF) — a forecast combination method in which the next-step point forecast is obtained by adjusting the original naive forecast with a movement prediction term.
+The weight assigned to this term is computed as the product of two in-sample quantities: a coefficient determined by the movement prediction accuracy and the mean absolute increment.
+The effectiveness of the proposed method was evaluated through experiments on eight real-world financial time series from the U.S.
+market.
+For all the experiments, the naive, naive with drift, integrated moving average of order (1,1) (IMA (1,1)), and linear regression serve as the baselines, with the RMSE, mean absolute error (MAE), mean absolute percentage error (MAPE) and symmetric mean absolute percentage error (sMAPE) used as evaluation metrics.
 This study makes three significant contributions to the literature:
-  • First, this study offers new insights into the predictability of random walk-like series, presenting a feasible approach to address this longstanding challenge in the field.
-When meaningful movement prediction is available, the point values of these series, which are traditionally viewed as unpredictable, exhibit conditional predictability.
-  • Second, the proposed method bridges movement prediction and point forecasting, which are two distinct tasks in time series forecasting.
-Although the use of binary classification to guide regression is well established, relying solely on the movement prediction of the target variable for its own point forecasting offers a fresh perspective.
-  • Third, this study introduces a clear boundary criterion, which explicitly states that the adjusted naïve forecast outperforms the original naïve forecast when the product of the actual optimal parameter values exceeds half the product of the estimated optimal parameter values.
-This contribution advances forecasting methodologies by clarifying the conditions required for improved performance.
+  • First, it proposes a novel forecast combination method that explicitly integrates movement predictions into point forecasting, uniquely linking these two distinct forecasting tasks.
+Although binary classification to guide regression is not new, exclusively using the target variable’s own movement prediction to inform its point forecast offers a fresh methodological perspective in forecasting.
+  • Second, it offers the first theoretical justification and mathematical analysis demonstrating how and why movement predictions alone can be exploited to generate point forecasts that outperform the naive forecast.
+This theoretical foundation highlights the informational value embedded in directional signals and distinguishes the proposed method from prior heuristic or empirical approaches.
+  • Third, it provides empirical insights into the conditional predictability of financial time series, demonstrating that series traditionally considered unpredictable can exhibit meaningful predictability when accurate movement prediction is available.
 The remainder of this paper is organized as follows.
-Section 2 introduces the details of the movement-prediction-adjusted naïve forecast, explains how to derive it and under what conditions it can outperform the original naïve forecast.
-Section 3 presents a simulation study that assesses the performance of movement-prediction-adjusted naïve forecasts via synthetic datasets in a controlled experimental setting.
-Section 4 discusses the key findings, implications, and limitations of the study.
+Section 2 introduces the details of the MPANF and explains how to derive it and under what conditions it can outperform the naive forecast.
+Section 3 presents empirical results on the performance of the MPANF via real-world financial time series.
+Section 4 discusses the implications of the findings and limitations of this study.
 Finally, section 5 concludes the paper and outlines future research directions.
 2 Method
-This section presents the details of the movement-prediction-adjusted naïve forecast.
-We first derive the general expression of the adjusted naïve forecast based on the definition of symmetric random walks and then explain how to optimize the parameters to improve the forecasting effectiveness and detail the procedure for generating forecasts on the out-of-sample set.
-2.1 General Formulation of the Movement-Prediction-Adjusted Naïve Forecast
-A symmetric random walk is a stochastic process in which the value at time step $t$ is the sum of the value at the previous time step $t-1$ and the value of an independent random increment ${{\epsilon}_{t}}$, which is generally assumed to follow a normal distribution with a mean of 0 and variance of $\sigma_{t}^{2}$ (De Gooijer et al.,, 2017).
-The formulation of symmetric random walks is as follows:
-${{y}_{t}}={{y}_{t-1}}+{{\epsilon}_{t}},{{\epsilon}_{t}}\sim\mathcal{N}(0,\sigma_{t}^{2}),$ (1)
-where ${y}_{t}$ is the point value at time step $t$ and ${y}_{t-1}$ is the point value at time step $t-1$.
-For random walks exhibiting homoscedasticity, the variance $\sigma_{t}^{2}$ remains constant over time, whereas for those exhibiting heteroscedasticity, the variance $\sigma_{t}^{2}$ varies over time.
-The symmetric nature of the random increment ${{\epsilon}_{t}}$ ensures that there is no inherent bias in the movement direction of the target variable and that upward and downward movements are equally likely (Ibe,, 2014).
-Furthermore, the random increment $\epsilon_{t}$ can be expressed as a product of two components: a nonnegative term $|\epsilon_{t}|$ representing the magnitude of the increment and a sign component ${D}_{t}$, which takes a value of $+1$ for upward movement and $-1$ for downward movement, indicating the direction of the increment between two adjacent points.
-This allows the process to be reformulated as follows:
+This section presents the details of the MPANF, which aims to outperform the naive forecast in point forecasting by incorporating the target variable’s movement prediction.
+We begin by deriving the general formulation of the MPANF.
+Subsequently, we optimize the necessary coefficient using the in-sample set and conduct uncertainty analysis for out-of-sample forecasts.
+2.1 General Formulation of the MPANF
+The value of a time series at time step $t$ can be expressed as the sum of its previous value at time step t1t-1 and an increment ${\epsilon}_{t}$}, which represents the value change between two adjacent observations.
+This relationship can be expressed as:
+${{y}_{t}}={{y}_{t-1}}+{{\epsilon}_{t}},$ (1)
+where ${y}_{t}$ is the point value at time step $t$ and ${y}_{t-1}$ is the point value at time step t1t-1.
+The increment $\epsilon_{t}$ at time step t can be expressed as a product of two terms: a nonnegative term $|\epsilon_{t}|$ representing the magnitude of the increment and a sign term ${D}_{t}$, which indicates the direction of the increment between two adjacent points and takes a value of +1+1 for upward movement and 1-1 for downward movement.
+This allows ${y}_{t}$ to be reformulated as follows:
 ${{y}_{t}}={{y}_{t-1}}+{{D}_{t}}\cdot|\epsilon_{t}|;|\epsilon_{t}|\geq 0$ (2)
@@
 ${{y}_{t}}={{y}_{t-1}}+{{D}_{t}}\cdot{\theta_{t}}\cdot\bar{\epsilon};{\theta_{t}}\geq 0,\bar{\epsilon}>0,$ (3)
-where $\bar{\epsilon}$ represents the mean absolute increment calculated based on the in-sample data, which can be interpreted as a base magnitude of the increments.
-The term $\theta_{t}$ is introduced as a nonnegative parameter that controls the magnitude of the increment at time step $t$.
+where $\bar{\epsilon}$ represents the mean absolute increment of the time series, which can be interpreted as a base magnitude of the increments.
+The term $\theta_{t}$ is introduced as a nonnegative coefficient that controls the magnitude of the increment at time step $t$.
 Eq.
-(3) implies that the point forecast of the target variable can be generated if ${{D}_{t}}$ can be predicted and both $\theta_{t}$ and $\bar{\epsilon}$ can be estimated.
-Consider a simplified scenario in which the magnitude of the increment between two adjacent points is fixed for future time steps; in this case, $\theta_{t}$ can be set to a constant value.
-Consequently, the point forecast for future time step $t$, denoted as $\hat{y}_{t}$, can be expressed as follows:
+(3) implies that the point forecast of the target variable can be generated if ${D}_{t}$} can be predicted and $\theta_{t}$ can be estimated.
+Consider a simple forecasting strategy in which the increments for each future time step are set to have a fixed magnitude; in other words, we set $\theta_{t}$ to be a constant value and derive the one-step-ahead point forecast at time step $t$, denoted as $\hat{y}_{t}$, as follows:
 ${{\hat{y}}_{t}}={{y}_{t-1}}+\hat{D}_{t}\cdot\theta\cdot\bar{\epsilon};\bar{\epsilon}>0,$ (4)
-where ${{\hat{D}}_{t}}$ is the movement prediction, $\theta$ is a fixed scalar, and $\theta\in\mathbb{Q}$.
-Suppose that the movement prediction has been provided, the point forecast generated via Eq.
-(4) can be regarded as an adjustment of the original naïve forecast through a weighted movement prediction term; hence, it is named the movement-prediction-adjusted naïve forecast.
-2.2 Parameter Optimization Using In-Sample Data
-According to Eq.
-(4), the weight of the movement term in the adjusted naïve forecast is the product of two parameters, $\theta$ and $\bar{\epsilon}$, both of which can be optimized based on the in-sample data.
-Because $\bar{\epsilon}$ is computed directly as the mean absolute increment of the in-sample series, only $\theta$ requires optimization.
-This optimization is typically achieved by minimizing the mean squared error (MSE) of the adjusted naïve forecast within the in-sample set.
-Given that the original naïve forecast’s MSE is constant for the in-sample set, minimizing the adjusted naïve forecast’s MSE is equivalent to maximizing their difference:
-$\min MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\equiv\max\left(MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\right)$ (5)
-Suppose the in-sample set has $N$ time steps, and movement predictions yield $n_{\text{in-sample}}^{\text{correct}}$ correct and $n_{\text{in-sample}}^{\text{incorrect}}$ incorrect directions; then, the directional accuracy of the movement prediction on the in-sample set can be expressed as follows:
+where ${\hat{D}}_{t}$} is the movement prediction, $\theta$ is a fixed scalar, and $\theta\in\mathbb{R}$.
+Therefore, the point forecast given by Eq.
+(4) is an adjusted naive forecast, where the adjustment is a movement prediction term weighted by the product of $\theta$ and $\bar{\epsilon}$.
+Hence, it is termed the movement-prediction-adjusted naive forecast (MPANF).
+2.2 Determination of the Optimal Coefficient $\theta$
+Since $\bar{\epsilon}$ in Eq.
+(4) is computed as the mean absolute increment of the in-sample set, only the coefficient $\theta$ needs to be optimized.
+Suppose that $\theta$ is optimized by minimizing the in-sample mean squared error (MSE) of the MPANF.
+Since the in-sample MSE of the original naive forecast is constant, this optimization is equivalent to maximizing the in-sample MSE difference between the original naive forecast and the MPANF:
+$\min MSE_{\text{in-sample}}^{\text{MPANF}}\equiv\max\left(MSE_{\text{in-sample}}^{\text{Naive}}-MSE_{\text{in-sample}}^{\text{MPANF}}\right),$ (5)
+where $MSE_{\text{in-sample}}^{\text{Naive}}$ and $MSE_{\text{in-sample}}^{\text{MPANF}}$ denote the in-sample MSE of the original naive forecast and that of the MPANF, respectively.
+Suppose the in-sample set has $N$ time steps, and the movement prediction yields $n_{\text{in-sample}}^{\text{correct}}$ correct and $n_{\text{in-sample}}^{\text{incorrect}}$ incorrect directions; then, the in-sample movement prediction accuracy, denoted as $ACC_{\text{in-sample}}$, can be expressed as follows:
 $ACC_{\text{in-sample}}=\frac{n_{\text{in-sample}}^{\text{correct}}}{n_{\text{in-sample}}^{\text{incorrect}}+n_{\text{in-sample}}^{\text{correct}}}$ (6)
-Accordingly, the MSE difference between two types of naïve forecasts is as follows:
-$\Delta MSE_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (7)
+Accordingly, the MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{in-sample}}$, is expressed as follows:
+$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{Naive}}-MSE_{\text{in-sample}}^{\text{MPANF}}$ (7)
 $=(4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}})\cdot\bar{\epsilon}_{\text{in-sample}}^{2},$
-where $\theta_{\text{in-sample}}$ represents the parameter $\theta$ in the adjusted naïve forecast for the in-sample set and $\bar{\epsilon}_{\text{in-sample}}$ denotes the mean of $|\epsilon_{t}|$ calculated over the in-sample set.
+where $\theta_{\text{in-sample}}$ denotes the coefficient $\theta$ used by the MPANF on the in-sample set and $\bar{\epsilon}_{\text{in-sample}}$ denotes the mean of $|\epsilon_{t}|$ calculated over the in-sample set.
 A detailed algebraic derivation of the $\Delta MSE_{\text{in-sample}}$ is provided in Appendix A.
-According to Eq. (7), the optimization reduces to maximizing the following function concerning parameter $\theta_{\text{in-sample}}$:
+According to Eq. (7), maximizing $\Delta MSE_{\text{in-sample}}$ can further be reduced to maximizing the following function of $\theta_{\text{in-sample}}$:
 $f(\theta_{\text{in-sample}})=4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}},$ (8)
-Differentiating Eq. (8) and setting the derivative to zero yields the optimal solution:
-Proposition 1 (Optimal parameter solution).
-The optimal parameter value $\theta_{\text{in-sample}}^{*}$ that maximizes the MSE difference between the adjusted naïve forecast and the original naïve forecast on the in-sample set is given by:
+Since Eq. (8) is a concave function, differentiating it with respect to $\theta_{\text{in-sample}}$ and setting the derivative to zero yields the optimal solution:
+Proposition 1 (Optimal coefficient).
+The optimal in-sample coefficient $\theta_{\text{in-sample}}^{*}$, which maximizes the in-sample MSE difference between the naive forecast and the MPANF, is given by:
 $\theta_{\text{in-sample}}^{*}=2\cdot ACC_{\text{in-sample}}-1$ (9)
-This result explicitly reveals that the optimal value of the parameter $\theta$ depends linearly on the directional accuracy of the movement predictions for a sequence with a fixed length.
-Since the second derivative of $f(\theta_{\text{in-sample}})$ is negative, this optimal solution is a maximum.
+This equation indicates that the optimal coefficient $\theta$ depends solely on the in-sample movement prediction accuracy and is independent of the specific movement predictions or the characteristics of the time series.
 Substituting Eq. (9) back into Eq. (7) yields:
-Proposition 2 (Performance guarantee).
-Under the optimal parameter setting using $\theta_{\text{in-sample}}^{*}$, the adjusted naïve forecast achieves a nonnegative improvement over the original naïve forecast in terms of the MSE on the in-sample set:
+Proposition 2 (Performance improvement).
+Under the optimal coefficient setting using $\theta_{\text{in-sample}}^{*}$, the MPANF achieves a nonnegative improvement over the naive forecast in terms of the in-sample MSE:
 $\max\Delta MSE_{\text{in-sample}}=\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\geq 0$ (10)
-Proposition 2 confirms that as the directional accuracy increases beyond 0.5, the value of $\theta_{\text{in-sample}}^{*}$ increases from 0, leading to progressively greater improvements over the original naïve forecast.
-At a directional accuracy of 1, the maximum achievable improvement is $\bar{\epsilon}_{\text{in-sample}}^{2}$.
-Thus, with the optimal parameter setting, the adjusted naïve forecast matches or surpasses the performance of the original naïve forecast within the in-sample dataset:
-$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\leq MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ (11)
-This result underlines the theoretical robustness of the proposed adjustment methodology.
-2.3 Uncertainty Analysis and Practical Generation of Out-of-Sample Forecasts
-We now examine the performance of the adjusted naïve forecast via out-of-sample data.
-Suppose that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are used as the parameters of the adjusted naïve forecast on the out-of-sample set, which contains $M$ time steps.
-Suppose that the movement predictions on this set yield $m_{\text{out-of-sample}}^{\text{correct}}$ correct and $m_{\text{out-of-sample}}^{\text{incorrect}}$ incorrect predictions; then, the directional accuracy of the movement prediction on the out-of-sample set can be expressed as:
+Proposition 2 confirms that as the directional accuracy increases beyond 0.5, the value of $\theta_{\text{in-sample}}^{*}$ increases from 0.0, leading to progressively greater improvements over the naive forecast.
+At a directional accuracy of 1.0, the maximum achievable improvement is $\bar{\epsilon}_{\text{in-sample}}^{2}$.
+Thus, with the optimal coefficient setting, the performance of the MPANF matches or surpasses that of the naive forecast on the in-sample set:
+$MSE_{\text{in-sample}}^{\text{MPANF}}\leq MSE_{\text{in-sample}}^{\text{Naive}}$ (11)
+2.3 Uncertainty Analysis of Out-of-Sample Forecasts
+We now examine the out-of-sample performance of the MPANF.
+Suppose that the movement predictions on the out-of-sample set with $M$ time steps yield $m_{\text{out-of-sample}}^{\text{correct}}$ correct and $m_{\text{out-of-sample}}^{\text{incorrect}}$ incorrect predictions; then, the out-of-sample movement prediction accuracy, denoted as $ACC_{\text{out-of-sample}}$, can be expressed as:
 $ACC_{\text{out-of-sample}}=\frac{m_{\text{out-of-sample}}^{\text{correct}}}{m_{\text{out-of-sample}}^{\text{incorrect}}+m_{\text{out-of-sample}}^{\text{correct}}}$ (12)
-Consequently, the MSE difference between two types of naïve forecasts on the out-of-sample is expressed as:
-$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (13)
+Suppose that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are used by the MPANF on the out-of-sample set; then, the out-of-sample MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{out-of-sample}}$, is expressed as:
+$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{Naive}}-MSE_{\text{out-of-sample}}^{\text{MPANF}}$ (13)
 $=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2},$
-where $\bar{\epsilon}_{\text{out-of-sample}}$ denotes the actual mean absolute increment on the out-of-sample set and $\theta_{\text{out-of-sample}}^{*}$ denotes the actual optimal value of $\theta$ for the out-of-sample set, which is derived according to Proposition 1:
+where $\bar{\epsilon}_{\text{out-of-sample}}$ denotes the out-of-sample mean absolute increment and $\theta_{\text{out-of-sample}}^{*}$ denotes the optimal out-of-sample coefficient $\theta$, which is derived according to Proposition 1:
 $\theta_{\text{out-of-sample}}^{*}=2\cdot ACC_{\text{out-of-sample}}-1.$ (14)
-A detailed algebraic expansions of the $\Delta MSE_{\text{out-of-sample}}$ is provided in Appendix B.
-For the adjusted naïve forecast to outperform the original naïve forecast, the $\Delta MSE_{\text{out-of-sample}}$ should be no less than zero.
+A detailed algebraic expansion of the $\Delta MSE_{\text{out-of-sample}}$ is provided in Appendix B.
+For the MPANF to outperform the naive forecast, the $\Delta MSE_{\text{out-of-sample}}$ should be no less than zero.
 Therefore, according to Eq.
@@
 $\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}\geq\frac{1}{2}\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}$ (15)
-Given that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are specific estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, re2spectively, inequality (15) can be generalized as:
-Proposition 3 (Out-of-sample performance criterion).
-The necessary condition for the adjusted naïve forecast to achieve an improvement over the original naïve forecast in terms of the MSE on the out-of-sample set is as follows:
+Given that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, respectively, inequality (15) can be generalized as:
+Proposition 3 (Out-of-sample performance criterion). The necessary condition for the MPANF to achieve an improvement over the naive forecast on the out-of-sample set in terms of the MSE is as follows:
 $\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}\geq\frac{1}{2}\cdot\hat{\theta}_{\text{out-of-sample}}\cdot\hat{\bar{\epsilon}}_{\text{out-of-sample}},$ (16)
-where $\hat{\theta}_{\text{out-of-sample}}$ represents the estimate of $\theta_{\text{out-of-sample}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ represents the estimate of $\bar{\epsilon}_{\text{out-of-sample}}$.
-Proposition 3 establishes a clear boundary criterion when an adjusted naïve forecast can outperform the original naïve forecast, specifically depending on whether the product of the actual optimal parameter values exceeds half the product of the estimated optimal parameter values.
-We can treat $\theta_{\text{out-of-sample}}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ as random variables, each following a distribution defined over nonnegative values.
-Accordingly, inequality (16) represents a probabilistic event, the likelihood of which depends on the realized values of these variables.
-Since $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ cannot be known in advance, the fulfillment of inequality (16) is uncertain.
-Larger parameter values have a lower likelihood of yielding improvements, although such improvements can be substantial, whereas smaller parameter values increase the likelihood of improvement, albeit potentially marginal or negligible, compared with the naïve forecast.
-Therefore, this boundary criterion highlights a critical trade-off between the likelihood of achieving improvements and the potential magnitude of these improvements.
-In addition to using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, a more conservative approach can be used to generate the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$:
-  1. 1.
-    Consider a data-generating process in which a sliding window, whose length is equal to that of the out-of-sample set, moves sequentially from the beginning to the end of the in-sample set.
-Each window can be treated as an "out-of-sample set" in the historical context.
-  2. 2.
-    Within each window, the directional accuracy is calculated and then translated into the corresponding value of $\theta$ via Eq.
-(9).
-As the window moves, a collection of $\theta$ values is generated, denoted as $\Theta_{\text{set}}$, whose bounds can be regarded as the empirical range of $\theta_{\text{out-of-sample}}^{*}$.
-  3. 3.
-    Similarly, a sliding window of the same length is applied to calculate a set of $\bar{\epsilon}$ values from the in-sample set, denoted as $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, providing an empirical value range of $\bar{\epsilon}_{\text{out-of-sample}}$.
-  4. 4.
-    Using $\min(\Theta_{\text{set}})$ and $\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})$ as conservative estimates for $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, the adjusted naïve forecast for time step $t$ in the out-of-sample set can be expressed as:
-    $\hat{y}_{t}=y_{t-1}+\hat{D}_{t}\cdot\min(\Theta_{\text{set}})\cdot\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})$ (17)
-Algorithm 1 describes the process of generating the adjusted naïve forecast on the entire out-of-sample set.
-Algorithm 1 Movement-Prediction-Adjusted Naïve Forecast on the Out-of-sample Set
-Inputs:
-$Y_{\text{in-sample}}$ //    Array of observations in the in-sample set
-$Y_{\text{out-of-sample}}$ //    Array of observations in the out-of-sample set
-$\hat{D}_{\text{in-sample}}$ //    Movement prediction on the in-sample set
-$\hat{D}_{\text{out-of-sample}}$ //    Movement prediction on the out-of-sample set
-$N$ //    Number of elements in the in-sample set
-$M$ //    Number of elements in the out-of-sample set
-Output:
-$P$ //    Adjusted naïve forecast on the out-of-sample set
-1. Initialize variables:
-$D_{\text{in-sample}}\leftarrow[]$ //    Actual movements on the in-sample set
-$P\leftarrow[]$
-$ACC_{\text{set}}\leftarrow[]$ //    Directional accuracy set
-$\Theta_{\text{set}}\leftarrow[]$ //    Theta set
-$R_{\text{in-sample}}\leftarrow[]$ //    First-order difference set
-$\boldsymbol{\bar{\epsilon}}_{\text{set}}\leftarrow[]$ //    Mean absolute increment set
-2. Calculate actual movements:
-for $t\leftarrow 1$ to $\text{length}(Y_{\text{in-sample}})$ do
-     $\text{diff}\leftarrow Y_{\text{in-sample}}[t]-Y_{\text{in-sample}}[t-1]$ //    First-order difference
-     $R_{\text{in-sample}}.\text{append}(\text{diff})$
-     if $\text{diff}>0$ then
-         $D_{\text{in-sample}}.\text{append}(1)$ //    Upward movement
-     else
-         $D_{\text{in-sample}}.\text{append}(-1)$ //    Downward or no movement
-     end if
-end for
-3. Calculate directional accuracies:
-for $t\leftarrow 1$ to $N-M+1$ do
+where $\hat{\theta}_{\text{out-of-sample}}$ denotes the estimate of $\theta_{\text{out-of-sample}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ denotes the estimate of $\bar{\epsilon}_{\text{out-of-sample}}$.
+Proposition 3 establishes a boundary criterion when the MPANF can outperform the naive forecast.
+We can treat $\theta_{\text{out-of-sample}}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ as two random variables, each following a distribution defined over nonnegative values.
+Accordingly, inequality (16) represents a probabilistic event whose likelihood depends on the realized values of these two variables.
+Smaller values of these two variables generally increase the likelihood of improvement.
+However, as shown in Eq.
+(13), the resulting gains may be marginal.
+The MPANF can realize its maximum potential improvement only when $\hat{\theta}_{\text{out-of-sample}}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ are close to their respective true values, $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$.
+In practice, the MPANF on the out-of-sample set can be derived via Algorithm 1.
+Algorithm 1 Movement-prediction–adjusted naive forecast on the out-of-sample set
+1:Inputs:
 […]
-     for $i\leftarrow t$ to $t+M-1$ do
-         if $D_{\text{in-sample}}[i]==\hat{D}_{\text{in-sample}}[i]$ then
 […]
-         end if
-     end for
-     $ACC_{\text{set}}.\text{append}(\text{Correct\_Predictions}/M)$
-end for
-4. Compute the movement weights:
-for each $\text{each }ACC\in ACC_{\text{set}}$ do
-     $\theta\leftarrow 2\cdot\text{ACC}-1$
-     $\Theta_{\text{set}}.\text{append}(\theta)$
-end for
-5. Compute the mean absolute increments:
-for $t\leftarrow 1$ to $N-M+1$ do
-     $\bar{\epsilon}\leftarrow\text{mean}(|R_{\text{in-sample}}[t:t+M-1]|)$
-     $\boldsymbol{\bar{\epsilon}}_{\text{set}}.\text{append}(\bar{\epsilon})$
-end for
-6. Forecast the out-of-sample set:
-$Y_{\text{last}}\leftarrow Y_{\text{in-sample}}[\text{end}]$ //    Start from the last element of the in-sample set
-for $i\leftarrow 1$ to $\text{length}(Y_{\text{out-of-sample}})-1$ do
-     $\Delta\leftarrow\min(\Theta_{\text{set}})\cdot\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})\cdot\hat{D}_{\text{out-of-sample}}[i]$ //    Compute increment
-     $Y_{\text{forecast}}\leftarrow Y_{\text{last}}+\Delta$ //    Compute the adjusted naïve forecast
-     $P.\text{append}(Y_{\text{forecast}})$ //    Store result
-     $Y_{\text{last}}\leftarrow Y_{\text{out-of-sample}}[i]$
-end for
-7. Return:
-return $P$
-3 Simulation
-This section presents a simulation study using four synthetic time series to evaluate whether the adjusted naïve forecast outperforms the original naïve forecast when movement predictions with different directional accuracy levels are provided.
-The performance of the proposed method under each directional accuracy level was evaluated based on the out-of-sample set of each time series.
+[…]
+[…]
+6:Output:
+7:$P$ $\triangleright$ Final point forecasts
+8:Initialize:
+[…]
+10:Step 1: Determine $\theta$ and $\bar{\epsilon}$ from in-sample data
+[…]
+[…]
+[…]
+[…]
+[…]
+16:Step 2: Generate forecasts on the out-of-sample set
+[…]
+18:for $i\leftarrow 1,\dots,\text{length}(Y_{\text{out-of-sample}})$ do
+19:  if $\hat{D}_{\text{out-of-sample}}[i]=1$ then
+20:   $\Delta\leftarrow\theta\cdot\bar{\epsilon}$ $\triangleright$ Upward
+21:  else
+22:   $\Delta\leftarrow-\theta\cdot\bar{\epsilon}$ $\triangleright$ Downward
+23:  end if
+24:  $\hat{Y}_{i}\leftarrow Y_{\text{last}}+\Delta$
+25:  $P.\text{append}(\hat{Y}_{i})$
+26:  $Y_{\text{last}}\leftarrow Y_{\text{out-of-sample}}[i]$
+27:end for
+28:return $P$
+3 Empirical Results
+This section presents the empirical evaluation of the MPANF using real-world data.
+It includes a description of the datasets, details of the experimental setup, and a performance comparison of the proposed method against several baselines.
 3.1 Datasets and Setup
-The dataset includes four synthetic time series, each of which represents a symmetric random walk with a certain type of variance, as summarized in Table 1.
-Each time series, consisting of 2,500 steps, was generated using a fixed random seed (seed = 1) in the Python environment, as shown in Figure 2.
-To maintain all values within the positive range, an initial positive offset of 10000 units was applied to each series.
-The last 500 data points were selected as the out-of-sample set, and the remaining data points were used as the in-sample set.
-In addition, Figure 3 shows the first-order differences of each time series to visually demonstrate the properties of the increments, and Figure 4 shows the histograms of the first-order differences for all time series, highlighting the symmetric nature of the time series.
-Table 1: Variance of synthetic random walks
-Variance Type Expression Parameter Range
-Constant variance $\sigma_{t}^{2}=\sigma_{0}^{2}$ $\sigma_{0}^{2}=1$
-Linear variance trend $\sigma_{t}^{2}=\frac{\sigma_{0}^{2}}{1+k\cdot t}$ $\sigma_{0}^{2}=1$, $k\in(0,10)$
-Cyclic variance $\sigma_{t}^{2}=\sigma_{0}^{2}(1+a\cdot\sin(\omega t))$ $\sigma_{0}^{2}=1$, $\omega=\frac{2\pi}{100}$, $a\in(0,10)$
-Randomly varying variance $\sigma_{t}^{2}=\sigma_{t-1}^{2}+\eta_{t},\eta_{t}\sim\mathcal{N}(0,\xi^{2})$ $\sigma_{0}^{2}=1$, $\xi^{2}\in(0,1000)$
-Figure 2: Synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
-Figure 3: First-order differences of four synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
-Figure 4: Histograms of first-order differences of four synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
-For each time series, the adjusted naïve forecast on the out-of-sample set was generated via Algorithm 1.
-The performance of each adjusted naïve forecast result was evaluated via four metrics: the RMSE, MAE, MAPE, and sMAPE.
+In the empirical analysis, we first generate movement predictions for each target series and then apply the MPANF to produce the final point forecasts.
+Since developing a complex movement prediction method is beyond the scope of this study, the movement predictions are obtained using a simple co-movement–based approach with a selected exogenous variable, as detailed in Appendix C.
+This two-step design reflects the nature of the MPANF as a second-stage method that relies on accurate and consistent directional inputs.
+Given the availability of data and the well-documented strong co-movement between the Financial Times Stock Exchange (FTSE) index and U.S.
+stock prices (Sarwar,, 2020; Gupta et al.,, 2024), together with the fact that the FTSE open price is recorded approximately 13 hours before the U.S.
+market close price on the same date, we selected eight U.S.
+financial time series as the target series, including Apple (AAPL), Boeing (BA), Google (GOOGL), Goldman Sachs (GS), Halliburton (HAL), Brent crude oil (OIL), Schlumberger (SLB), and Tesla (TSLA), based on their strong predictive alignment with FTSE movements.
+Accordingly, the FTSE open price is used as the exogenous variable for generating movement predictions.
+Historical data for the FTSE open price and the close prices of the eight target series were obtained from Yahoo Finance, covering the period from October 8, 2014, to September 13, 2024.
+The FTSE data were aligned with each target series on a day-by-day basis, with extra observations removed and missing values imputed via forward filling (Junninen et al.,, 2004).
+After alignment, all series were standardized to a uniform length of 2,500 observations and then split into in-sample and out-of-sample sets at a 1:1 ratio, ensuring sufficient data for both model estimation and evaluation.The close prices of the eight target variables and the FTSE open price are illustrated in Figures 2 and 3, respectively, with a brief description provided in Table 1.
+The in-sample movement prediction accuracies for the eight target series all exceeded 0.55, indicating that the directional signals derived from the FTSE open price were meaningful.
+Figure 2: Time series of eight target variables.
+Figure 3: Open price of FTSE.
+Table 1: Statistics of nine financial time series.
+Time series Count Mean Std. Dev. Minimum Median Maximum $\mathbf{ACC}_{\text{in-sample}}$ $\mathbf{\bar{\epsilon}}_{\text{in-sample}}$
+AAPL (Close price) 2500 88.46 62.00 20.72 54.16 234.55 0.5536 0.3842
+BA (Close price) 2500 211.21 80.90 95.01 195.92 430.30 0.5568 2.5867
+GOOGL (Close price) 2500 78.45 41.52 24.79 61.44 190.93 0.5504 0.4624
+GS (Close price) 2500 239.35 89.70 117.58 202.27 510.25 0.5504 1.9030
+HAL (Close price) 2500 30.99 9.61 4.32 33.30 151.09 0.5680 0.5252
+OIL (Close price) 2500 66.43 18.59 19.33 65.20 127.98 0.5576 0.9067
+SLB (Close price) 2500 45.94 15.45 11.02 49.49 74.91 0.5536 0.6596
+TSLA (Close price) 2500 107.38 109.43 9.58 23.72 409.97 0.5536 0.3534
+FTSE (Open price) 2500 7088.28 589.13 4993.90 7210.00 8445.80
+Since many sophisticated models, such as machine learning or deep learning-based models, struggle to provide significant improvements over naive forecasts in financial time series forecasting, simple but strong baselines, such as naive, naive with drift, IMA (1,1), and linear regression, are selected for performance comparison Rogers, (1995); Jadevicius and Huston, (2015); Franses, (2020).
+In particular, the linear regression model, which acts as a meta-learner to combine the naive forecast and the movement prediction, uses only the current observed value and the predicted movement as predictors.
+The performance of each method was evaluated via four metrics: the RMSE, MAE, MAPE, and sMAPE.
 The RMSE, derived from the square root of the MSE, emphasizes larger errors, which makes it suitable for identifying significant deviations in forecasts.
@@
 where $N$ denotes the size of the out-of-sample set, ${\hat{y}}_{i}$} denotes the predicted value, and ${y}_{i}$} denotes the actual value.
-To statistically test the performance of the adjusted naïve forecast, for each synthetic time series at a specified directional accuracy level, the process to generate the adjusted naïve forecast was repeated 100 times.
-Each repetition uses a unique movement prediction sequence at the same directional accuracy level, which is generated by randomly replacing a subset of actual movements in the out-of-sample set with incorrect movements.
-The directional accuracy level varies in two ways: initially, it changes from 0.50 to 0.56 in increments of 0.01, and then it changes from 0.55 to 1.00 in larger increments of 0.05.
-This approach allows for finer granularity in the lower accuracy range, where small changes in the directional accuracy can have a significant effect on the performance of the method, while using larger steps at higher accuracy levels balances the computational efficiency.
-Consequently, for each synthetic time series, 100 independent values were calculated for each metric at each accuracy level.
 3.2 Results and Analysis
-Because the directional accuracy for the out-of-sample set is predefined in each simulation process, the value of $\theta_{\text{out-of-sample}}^{*}$ can be derived directly via Eq.
-(14), and only the value of $\bar{\epsilon}_{\text{out-of-sample}}$ needs to be estimated.
-Figure 5 illustrates the distribution of the values in the $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, generated based on the sliding windows across the first-order differences of the in-sample set.
-The vertical reference line in Figure 5 represents the actual value of $\bar{\epsilon}_{\text{out-of-sample}}$.
-In the four cases, the minimum value of $\boldsymbol{\bar{\epsilon}}_{\text{set}}$ is generally an underestimate of $\bar{\epsilon}_{\text{out-of-sample}}$, except for the second case, where the synthetic time series is a random walk with a decreasing trend of linear variance.
-In general, the estimated value remains close to the true value.
-Under the assumption that the directional accuracy is consistent across the in-sample and out-of-sample sets, this estimation ensures that inequality (16) is satisfied.
-Figure 5: Comparison of the true $\bar{\epsilon}_{\text{out-of-sample}}$ with $\boldsymbol{\bar{\epsilon}}_{\text{set}}$: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
-Furthermore, comparisons of the two types of naïve forecasts based on the four evaluation metrics at low accuracy levels are illustrated in Figures 6, 7, 8, and 9.
-Each box plot in each figure represents the distribution of 100 independent values for one metric at a specific directional accuracy level, with the metric value of the original naïve forecast shown as a reference line.
-According to Figure 6, at low accuracy levels, such as 0.51, the RMSE of the adjusted naïve forecast is comparable to or slightly better than that of the original naïve forecast, as shown by the proximity of the median of each box plot to the reference line.
-As the accuracy increased beyond 0.52, most of the RMSE values fell below the baseline.
-Figure 6: RMSEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
-Figure 7: MAEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
-Figure 8: MAPEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
-Figure 9: sMAPEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
-Because the adjusted naïve forecast consistently outperforms the original naïve forecast at low accuracy levels in terms of the MAE, MAPE, and sMAPE, the Wilcoxon signed rank test was conducted to specifically evaluate whether the median differences between the RMSE values of the adjusted naïve forecast and those of the original naïve forecast significantly deviated from zero for accuracy levels ranging from 0.50–0.56 (Li and Johnson,, 2014).
-As summarized in Table 2, at a directional accuracy level of 0.51, the adjusted naïve forecast shows mixed results with respect to statistical significance.
-Although significant improvements were observed over the original naïve forecast for the Type 4 series (randomly varying variance, p = 0.002), the results for the Type 1 (constant variance, p = 0.915), Type 2 (linear variance trend, p = 0.781), and Type 3 (cyclic variance, p = 0.812) series were not statistically significant, indicating that the adjusted naïve forecast under an accuracy level of 0.51 may not consistently produce meaningful improvements across all types of time series.
-As the accuracy increased beyond 0.51, significant reductions in the RMSE (p < 0.001) were observed for all types of time series, thereby demonstrating the reliability and effectiveness of the adjusted naïve forecast at higher accuracy levels.
-Table 2: RMSE results of the adjusted naïve forecast across directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
-Prediction Type 1 Type 2 Type 3 Type 4
-RMSE Range p-value RMSE Range p-value RMSE Range p-value RMSE Range p-value
-Original Naïve 0.961 - 0.00762 - 1.355 - 21.034 -
-Acc = 0.50 0.961 - 0.00762 - 1.355 - 21.034 -
-Acc = 0.51 0.960 $\pm$ 0.0004 0.915 0.00762 $\pm$ 0.000003 0.781 1.355 $\pm$ 0.0005 0.812 21.032 $\pm$ 0.0053 0.002
-Acc = 0.52 0.960 $\pm$ 0.0008 <0.001 0.00762 $\pm$ 0.000007 <0.001 1.355 $\pm$ 0.0008 <0.001 21.024 $\pm$ 0.0102 <0.001
-Acc = 0.53 0.959 $\pm$ 0.0012 <0.001 0.00761 $\pm$ 0.000010 <0.001 1.354 $\pm$ 0.0015 <0.001 21.015 $\pm$ 0.0144 <0.001
-Acc = 0.54 0.959 $\pm$ 0.0017 <0.001 0.00761 $\pm$ 0.000014 <0.001 1.353 $\pm$ 0.0020 <0.001 20.995 $\pm$ 0.0199 <0.001
-Acc = 0.55 0.957 $\pm$ 0.0020 <0.001 0.00760 $\pm$ 0.000020 <0.001 1.352 $\pm$ 0.0028 <0.001 20.979 $\pm$ 0.0244 <0.001
-Acc = 0.56 0.956 $\pm$ 0.0023 <0.001 0.00758 $\pm$ 0.000021 <0.001 1.351 $\pm$ 0.0032 <0.001 20.957 $\pm$ 0.0278 <0.001
-According to Figures 10, 11, 12, and 13, at higher accuracy levels ranging from 0.55–1.00, the adjusted naïve forecast consistently outperforms the original naïve forecast in terms of the four metrics.
-In particular, the improvement over the original naïve forecast becomes substantial as the directional accuracy improves.
-This finding demonstrates that the proposed method can effectively translate highly accurate movement predictions into improved point forecasts in the context of forecasting random walk-like time series.
-However, the improvement over the original naïve forecast has a limit.
-The proposed method allows a random walk-like series to be predicted to a certain extent, but the predicted value cannot be infinitely close to the actual value.
-Figure 10: RMSEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
-Figure 11: MAEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
-Figure 12: MAPEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
-Figure 13: sMAPEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
-In general, when the directional accuracy was 0.5, the adjusted naïve forecast did not provide an improvement over the original naïve forecast.
-As the directional accuracy increases, the performance of the adjusted naïve forecast improves.
-The simulation results also demonstrate that the effectiveness of the proposed method is unaffected by the variance type of the synthetic time series, as evidenced by the consistent performance of the adjusted naïve forecast across all synthetic series.
-This result is consistent with the theoretical analysis of the proposed method, which highlights its robustness and potential applicability to a wide range of time series that exhibit symmetric random walk characteristics, including those with homoscedasticity as well as high-volatility scenarios.
+The out-of-sample forecasting results for the eight financial time series are illustrated in Figure 4.
+Given that each out-of-sample set consists of 1250 time steps and that predictions from different methods visually overlap to a large extent, only the first five predicted points from each set are displayed for clarity.
+Small differences between the MPANF and the other forecasting methods are observed because the overall in-sample movement prediction accuracy for each time series is just slightly above 0.55, leading to a small coefficient $\theta$ (approximately 0.1) and a modest adjustment of the naive forecast at each time step.
+The naive forecast, naive forecast with drift, and IMA(1,1) forecasts yield similar predictions, reinforcing their suitability as baseline methods for time series forecasting.
+Moreover, because the linear regression model closely mirrors the structure of the MPANF, its performance is similarly dependent on movement prediction accuracy and therefore remains close to that of the MPANF.
+Figure 4: Out-of-sample forecasts by different methods.
+Tables 2, 3, 4, and 5 present performance comparisons between the MPANF and the baseline methods for eight out-of-sample sets, reported in terms of the RMSE, MAE, MAPE, and sMAPE, respectively.
+Notably, the MPANF consistently outperforms the naive forecast, with the only exceptions being the MAPE and sMAPE for the GOOGL dataset.
+Specifically, the MPANF achieves the lowest RMSE in 4 out of 8 cases, the lowest MAE in 6 out of 8 cases, the lowest MAPE in 6 out of 8 cases, and the lowest sMAPE in 5 out of 8 cases, demonstrating its overall superiority over the baseline methods.
+Table 2: RMSE comparison across forecasting methods for selected stocks
+Stock Naive Naive + Drift IMA(1,1) LR MPANF
+AAPL 2.5937 2.5925 2.5979 2.5942 2.5931
+BA 5.9080 5.9187 5.9166 5.8932 5.8829
+GOOGL 2.2093 2.2085 2.2108 2.2114 2.2088
+GS 5.3671 5.3660 5.3727 5.4039 5.3485
+HAL 0.7334 0.7341 0.7332 0.7261 0.7271
+OIL 1.8909 1.8912 1.8976 1.8940 1.8864
+SLB 0.9401 0.9410 0.9401 0.9331 0.9343
+TSLA 8.1524 8.1525 8.1677 8.3351 8.1515
+Table 3: MAE comparison across forecasting methods for selected stocks
+Stock Naive Naive + Drift IMA(1,1) LR MPANF
+AAPL 1.8950 1.8935 1.8975 1.8958 1.8943
+BA 3.9691 3.9821 3.9750 3.9657 3.9533
+GOOGL 1.5680 1.5660 1.5727 1.5746 1.5674
+GS 3.9743 3.9737 3.9811 4.0021 3.9534
+HAL 0.5423 0.5429 0.5439 0.5376 0.5368
+OIL 1.2959 1.2981 1.3000 1.3073 1.2903
+SLB 0.6901 0.6910 0.6907 0.6852 0.6843
+TSLA 5.5732 5.5733 5.5775 5.8155 5.5722
+Table 4: MAPE (%) comparison across forecasting methods for selected stocks
+Stock Naive Naive + Drift IMA(1,1) LR MPANF
+AAPL 1.4081 1.4069 1.4065 1.4079 1.4074
+BA 2.0601 2.0675 2.0633 2.0585 2.0514
+GOOGL 1.4357 1.4340 1.4380 1.4406 1.4359
+GS 1.4037 1.4036 1.4046 1.4074 1.3967
+HAL 2.3517 2.3532 2.3667 2.3277 2.3216
+OIL 1.8379 1.8409 1.8430 1.8456 1.8296
+SLB 2.1522 2.1532 2.1574 2.1355 2.1291
+TSLA 2.9511 2.9513 2.9549 3.0534 2.9509
+Table 5: sMAPE (%) comparison across forecasting methods for selected stocks
+Stock Naive Naive + Drift IMA(1,1) LR MPANF
+AAPL 1.4088 1.4072 1.4072 1.4088 1.4081
+BA 2.0542 2.0591 2.0591 2.0500 2.0457
+GOOGL 1.4352 1.4331 1.4375 1.4410 1.4354
+GS 1.4035 1.4032 1.4045 1.4090 1.3966
+HAL 2.3377 2.3427 2.3537 2.3051 2.3084
+OIL 1.8290 1.8328 1.8336 1.8377 1.8210
+SLB 2.1465 2.1503 2.1521 2.1260 2.1240
+TSLA 2.9518 2.9520 2.9554 3.0767 2.9515
+In addition, baseline methods such as naive with drift, IMA(1,1), or linear regression occasionally yield lower RMSE values (as observed for AAPL, GOOGL, HAL, and OIL), lower MAE values (AAPL, GOOGL), lower MAPE values (AAPL, GOOGL), and lower sMAPE values (AAPL, GOOGL, and HAL).
+However, the performances of these methods are inconsistent; they frequently fall short of the naive forecast in other cases.
+Notably, although linear regression also uses movement prediction as an input feature and may occasionally outperform the naive forecast, it lacks a theoretical explanation for such improvement.
+Moreover, its performance generally lags behind that of the MPANF.
+This performance gap arises because linear regression is more susceptible to overfitting.
+While it closely fits the in-sample data, its out-of-sample performance can deteriorate significantly if the movement prediction accuracy decreases.
+In contrast, the coefficient determination in the MPANF is guided by theoretically derived optimal conditions rather than by data-driven minimization of in-sample errors.
+This design inherently introduces regularization, effectively mitigating overfitting and enhancing the method’s generalization ability.
+Additionally, Figures 5, 6, 7, and 8 present the performance of each forecasting method across the four evaluation metrics.
+Overall, the MPANF generates more accurate and robust forecasts than the baseline methods do and consistently improves upon the naive forecast across diverse financial forecasting scenarios.
+These empirical results are consistent with the theoretical analysis, underscoring the method’s broad applicability to financial time series forecasting, including series characterized by heteroscedasticity and high volatility.
+Figure 5: Comparison of the RMSEs of different methods.
+Figure 6: Comparison of the MAEs of different methods.
+Figure 7: Comparison of the MAPEs of different methods.
+Figure 8: Comparison of the sMAPEs of different methods.
 4 Discussion
-This section discusses the implications of the findings and potential limitations of this study.
-First, this study demonstrates an indirect regression approach that leverages movement predictions in conjunction with naïve forecasts for time series exhibiting random walk-like behavior.
-The theoretical analysis and simulation results suggest that accurate movement prediction can be systematically translated into improved point forecasts.
-This highlights the potential value of incorporating directional information into simple baseline models and suggests that enhancing movement prediction may, in some contexts, be more effective than developing increasingly complex direct regression models.
-Second, in many real-world forecasting scenarios, decision-making often involves integrating predictions with heterogeneous data types, such as combining binary classifications and continuous regressions.
-However, such integration can be technically challenging.
-The proposed method provides a practical means of converting binary movement predictions into continuous point forecasts for random walk-like time series, thereby enabling more informed decision-making in applied forecasting contexts.
-Third, the goal of this study is to examine the theoretical possibility of improving forecasts for random walk-like time series by incorporating directional signals rather than disputing the broader validity of the naïve forecast as the optimal benchmark.
-While the simulation implies that even weak directional signals can be exploited to enhance point forecasts, these improvements can be modest and context dependent.
-The adjusted naïve forecast serves as a second-stage method only after movement prediction is provided.
-Moreover, such improvements do not necessarily translate into economic gains, particularly when real-world constraints such as transaction costs, slippage, and market frictions are considered.
-Finally, because the focus of this study is on the theoretical impact of directional information rather than on modeling movement predictions themselves, we rely on controlled simulations rather than empirical data.
-The effectiveness of the proposed adjustment depends on the consistency of the directional accuracy across different segments of the series.
-In practice, movement prediction models susceptible to overfitting may suffer from deteriorated performance on out-of-sample data, thereby limiting the actual benefit of incorporating such predictions into the original naïve forecast.
-5 Conclusion
-This study introduces a movement-prediction-adjusted naïve forecast, which is a linear combination of movement prediction and the original naïve forecast, in the challenging context of forecasting time series that exhibit symmetric random walk characteristics.
-The weight of the movement term is determined by two parameters derived from the in-sample data: one based on the directional accuracy of the movement prediction and the other on the mean absolute increment of the target series.
-The theoretical analysis of the proposed method explains how accurate movement prediction can be exploited to enhance point forecasting for a random walk-like series.
-It also defines a boundary criterion stating that the adjusted naïve forecast outperforms the original naïve forecast when the product of the actual optimal parameter values exceeds half the product of their estimated values.
-Furthermore, the simulation results demonstrate that the adjusted naïve forecast can achieve statistically significant improvements over the original naïve forecast, even with directional accuracies slightly greater than 0.50.
-These findings highlight the significant potential of using movement predictions to address the inherent challenges of point forecasting in a random walk-like time series.
-There are several avenues for further research on the foundation of this study.
-First, while this research focuses on time series with symmetric random walk characteristics, future work could explore applying this movement prediction-based method to time series with nonsymmetric random walk properties, potentially broadening its applicability.
-Second, it is worth exploring whether the proposed forecasting method can be generalized beyond the one-dimensional case, as many real-world variables exhibit random walk behavior in a high-dimensional space.
-Moreover, the proposed method uses a fixed increment for simplicity.
-However, adapting this increment based on the availability of new data may provide more dynamic and accurate forecasts.
-Investigating how the increments can be adapted and whether such adaptive increments can further refine the forecasting accuracy represents a promising direction for future research.
+This section discusses the implications of the findings and the potential limitations of the study.
+First, the results demonstrate the effectiveness of a second-stage forecasting method that combines movement predictions with naive forecasts for financial time series forecasting.
+Both the theoretical analysis and the empirical evidence indicate that accurate movement prediction can be systematically translated into improved point forecasts.
+This underscores the value of incorporating directional information into a naive forecast and suggests that, in some contexts, the indirect forecasting approach may be more effective than developing complex regression models.
+Second, in many real-world forecasting applications, decision-making requires integrating predictions from heterogeneous data types, such as binary classifications and continuous regressions.
+Such integration can be technically challenging.
+The proposed method offers a practical solution by converting binary movement predictions into continuous point forecasts, thereby facilitating more informed decisions in forecasting scenarios.
+Third, while the empirical results show that even relatively weak directional signals can enhance point forecasts, the magnitude of such improvements is often modest and context dependent.
+Moreover, these gains do not necessarily translate into economic benefits, especially when real-world factors such as transaction costs, slippage, and market frictions are taken into account.
+Finally, although the MPANF imposes no restrictions on the choice of movement prediction models, its effectiveness depends critically on the consistency and reliability of directional accuracy across different segments of the dataset.
+In practice, predicting the movements of financial time series remains challenging, and models prone to overfitting may suffer substantial degradation in out-of-sample performance.
+Such deterioration can reduce, or even eliminate, the practical advantages of incorporating movement predictions into the naive forecast.
+In other words, the applicability of the MPANF is conditional on the availability of a meaningful and robust movement prediction for the target variable.
+5 Conclusions
+This study introduces the MPANF, a forecast combination method that adjusts the naive forecast using the movement prediction of the target variable.
+The approach is particularly effective for time series in which surpassing the naive forecast is typically difficult but accurate movement prediction is available.
+The weight assigned to the movement term is calculated as the product of a coefficient derived from the in-sample movement prediction accuracy and the in-sample mean absolute increment.
+Theoretical analysis explains why this combination can yield point forecasts that are superior to the naive forecast.
+Empirical evaluation on eight financial time series confirms the MPANF’s ability to achieve robust improvements over the naive forecast, even when the directional accuracy is only modestly above 0.55.
+Together, these findings underscore the potential of leveraging movement predictions to enhance point forecasts in challenging forecasting scenarios.
+Several avenues for future research emerge from this work.
+First, because the present evaluation was conducted under modest directional accuracy, developing more advanced movement prediction models and assessing MPANF performance at higher accuracy levels could enhance its practical value in real-world trading.
+Second, applying MPANF to other time series domains—particularly those where the naive forecast is hard to beat yet movement prediction is feasible—could broaden its applicability.
+Third, extending the MPANF beyond the univariate case would address the many forecasting tasks that arise in high-dimensional settings.
+Finally, while this study employs a fixed adjustment for simplicity, exploring adaptive adjustment schemes may further improve the forecasting accuracy.
 Acknowledgments
 This study did not receive any specific grants from funding agencies in the public, commercial, or not-for-profit sectors.
-Appendix A. Algebraic Derivation of $\Delta MSE_{\text{in-sample}}$
-According to Eq. (2), which presents the actual value of the target variable, the MSE expression for the original naïve forecast can be expressed as:
-$MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (A1)
+Appendix A. Algebraic Derivation of $\Delta MSE_{\textnormal{in-sample}}$
+According to Eq. (2), the in-sample MSE expression for the naive forecast, denoted as $MSE_{\text{in-sample}}^{\text{Naive}}$, can be expressed as:
+$MSE_{\text{in-sample}}^{\text{Naive}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (A1)
 $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t-1}+D_{t}\cdot|\epsilon_{t}|-y_{t-1}\right)^{2}\right)$
 $=\frac{1}{N}\left(\sum_{t\in\text{correct}}\ |\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}\ |\epsilon_{t}|^{2}\right)$
-Similarly, by substituting Eq. (2) for actual values and Eq. (4) for predicted values, the MSE of the adjusted naïve forecast can be expressed as follows:
-$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$ (A2)
+By substituting Eq. (2) into $y_{t}$ and Eq. (4) into $y_{t-1}$, the in-sample MSE of the MPANF, denoted as $MSE_{\text{in-sample}}^{\text{MPANF}}$, can be expressed as follows:
+$MSE_{\text{in-sample}}^{\text{MPANF}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$ (A2)
 $=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(y_{t}-\hat{y}_{t}\right)^{2}+\sum_{t\in\text{incorrect}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
@@
 $\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right),$
-Hence, the MSE difference between the adjusted naïve forecast and the original naïve forecast on the in-sample set can be derived via Eq. (A1) and Eq. (A2):
-$\Delta\text{MSE}_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (A3)
+Hence, the in-sample MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{in-sample}}$, can be derived via Eq. (A1) and Eq. (A2):
+$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{Naive}}-MSE_{\text{in-sample}}^{\text{MPANF}}$ (A3)
 $=\frac{1}{N}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
@@
 $\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
-Based on the law of large numbers (LLN), when both $n_{\text{in-sample}}^{\text{correct}}$ and $n_{\text{in-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
+Based on the law of large numbers (LLN), under idealized conditions where both $n_{\text{in-sample}}^{\text{correct}}$ and $n_{\text{in-sample}}^{\text{incorrect}}$ are sufficiently large, and the two sets of increments corresponding to correct and incorrect movement predictions are unbiased and representative of the overall increment set, the following approximations hold:
 $\sum_{t\in\text{correct}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}}$ (A4)
 Therefore, Eq. (A3) can be simplified as follows::
-$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (A5)
+$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{Naive}}-MSE_{\text{in-sample}}^{\text{MPANF}}$ (A5)
 $\approx\frac{1}{N}\left(2\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}})-n_{\text{in-sample}}^{\text{correct}}\cdot\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right$.
@@
 $=(4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}})\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
-Appendix B. Algebraic Derivation of $\Delta MSE_{\text{out-of-sample}}$
-When using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as parameters for the adjusted naïve forecast on the out-of-sample set, the difference in MSE between the adjusted and original naïve forecasts on the out-of-sample set can be expressed as follows:
-$\Delta\text{MSE}_{\text{out-of-sample}}$ $=\text{MSE}_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (B1)
+Appendix B. Algebraic Derivation of $\Delta MSE_{\textnormal{out-of-sample}}$
+When $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are used by the MPANF on the out-of-sample set, the out-of-sample MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{out-of-sample}}$, can be expressed as follows:
+$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{Naive}}-MSE_{\text{out-of-sample}}^{\text{MPANF}}$ (B1)
 $=\frac{1}{M}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
@@
 $\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
-According to the LLN, when both $m_{\text{out-of-sample}}^{\text{correct}}$ and $m_{\text{out-of-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
+According to the LLN, under idealized conditions where both $m_{\text{in-sample}}^{\text{correct}}$ and $m_{\text{in-sample}}^{\text{incorrect}}$ are sufficiently large, and the two sets of increments corresponding to correct and incorrect movement predictions are unbiased and representative of the overall increment set, the following approximations hold:
 $\sum_{t\in\text{correct}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}}$ (B2)
 Therefore, Eq. (B1) can further be expressed as:
-$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (B3)
+$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{Naive}}-MSE_{\text{out-of-sample}}^{\text{MPANF}}$ (B3)
 $\approx\frac{1}{M}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}})-m_{\text{out-of-sample}}^{\text{correct}}\cdot\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right$.
@@
 $=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
+Appendix C. Movement Prediction with an Exogenous Variable
+A simple but effective method for generating movement prediction involves leveraging the co-movement between the target variable and an exogenous variable.
+The selection of the exogenous variable is typically guided by domain knowledge, and the movement prediction of the target variable, denoted as $\hat{D}_{t}^{\text{target}}$, can be simply obtained as:
+$\hat{D}_{t}^{\text{target}}=\hat{D}_{t}^{\text{exo}},$ (C1)
+where $\hat{D}_{t}^{\text{exo}}$ represents the movement prediction of the exogenous variable at time step $t$.
+In particular, if the value of the exogenous variable is recorded sufficiently ahead of the target variable at the same time step, such as in financial markets operating in different time zones, the actual movement of the exogenous variable can directly serve as the movement prediction for the target variable:
+$\hat{D}_{t}^{\text{target}}=D_{t}^{\text{exo}}=\begin{cases}1,&\text{if }y_{t}^{\text{exo}}>y_{t-1}^{\text{exo}};\\ -1,&\text{if }y_{t}^{\text{exo}}\leq y_{t-1}^{\text{exo}};\end{cases},$ (C2)
+where $D_{t}^{\text{exo}}$ is the actual movement of the target variable at time step $t$.
+$y_{t}^{\text{exo}}$ and $y_{t-1}^{\text{exo}}$ denotes the actual values of the exogenous variable at time steps $t$ and t1t-1, respectively.
+The in-sample movement prediction accuracy of the target variable, denoted as $ACC_{\text{in-sample}}$, can be obtained as follows:
+$ACC_{\text{in-sample}}=\frac{1}{n}\sum\limits_{i=1}^{n}{1(D_{i}^{\text{exo}}=D_{i}^{\text{target}})},$ (C3)
+where $n$ is the total number of time steps in the in-sample set and $D_{i}^{\text{exo}}$ and $D_{i}^{\text{target}}$ are the $i\text{th}$ actual movements of the exogenous and target variables within the in-sample set, respectively.
+$1(D_{i}^{\text{exo}}=D_{i}^{\text{target}})$ is an indicator function that equals 1 if the movement prediction of the exogenous variable matches the movement prediction of the target variable and 0 otherwise.
+Only when the movement prediction accuracy is greater than a predefined threshold (e.g., 0.55) is the movement prediction considered meaningful and suitable for using the MPANF at the second stage.